### v4

# DAIRstega: Dynamically Allocated Interval-Based Generative Linguistic Steganography with Roulette Wheel

Abstract
Linguistic steganography (LS) tasks aim to generate steganographic text (stego) based on secret. Only authorized receivers can perceive and extract secrets, thereby protecting privacy. However, existing generative LS schemes often do not consider the conditional probability of tokens in the candidate pool, and allocate one or the same number of codings to all tokens. The tokens with lower probabilities are selected to embed secrets that will affect the quality of stegos. As a result, the stegos are easy to perceive and detect. This paper proposes the LS scheme based on dynamically allocated intervals, called DAIRstega. DAIRstega uses the idea of the roulette wheel and takes the conditional probabilities of tokens as the main basis for allocating the roulette area (i.e., the interval length). Thus, the token with a larger conditional probability is allocated more. The secret will be more likely to select the tokens with larger probabilities. In the allocation process, we design some functions between probability and allocated interval length. Based on the invisible characteristics of LS, we give three constraints that need to be met to design the function. To simplify the form, the expression of the allocation function is condensed. Furthermore, DAIRstega can receive additional instruction and controllably generate stegos. Rich experiments show that the proposed embedding way and DAIRstega perform superior to the existing ways and LS schemes, which shows strong perceptual, statistical, and semantic concealment and anti-steganalysis ability. This scheme can also generate high-quality longer stegos, improving the deficiencies in this task. The experiment also verified that DAIRstega can be used as a secure watermarking scheme, providing some ideas for its development.
1 Introduction
Steganography, a technology of information hiding (Shannon 1949), embeds secrets within media such as texts (Yang et al. 2023, 2021) and images, obtaining steganographic media that is sensory indistinguishable from normal media. Only authorized receivers can perceive whether the media is steganographic and accurately extracts the secret. Benefiting from the ability of text to be transmitted losslessly over public channels, research on Linguistic Steganography (LS) has explosive growth (Lu et al. 2023; Wang et al. 2023a; Yang et al. 2023). To increase the anti-detection ability of steganographic text (stego), the focus of work has shifted from modified LS (Huo and Xiao 2016) represented by synonym replacement (Kim, Zaiane, and Goebel 2010) to generative LS based on secret (Li et al. 2021; Yang et al. 2021).
The concealment of stego determines the success of covert communication to a great extent. Depending on the constraints, the concealment is manifested in three aspects: perceptual, statistical, and semantic. “Perceptual concealment” focuses on ensuring stegos with complement and fluency. “Statistical concealment” requires the distribution of stegos to be close to that of the natural texts (covers). “Semantic concealment” aims to stego that is coherent and semantically controllable. The concealment of stego mainly depends on the design of the steganography embedding way. An excellent embedding way ensures that the conditional probability (CP) distribution of the text is not changed as much as possible while embedding the secret. In this way, stegos are difficult to perceive and detect by unauthorized receivers. Most existing LS schemes have made efforts in this regard. Yang et al. (Yang et al. 2019) and Ding et al. (Ding et al. 2023a) encoded each token in the candidate pool with Fixed length coding (FLC) and only use secret information to determine the next token to achieve the purpose of embedding. Yang et al. (Yang et al. 2021) and Zhou et al. (Zhou et al. 2021) designed the Huffman coding-based (HC) embedding, which generates stegos with good completeness and fluentness, improving perceptual concealment. Zhang et al. (Zhang et al. 2021) and Wang et al. (Wang et al. 2023a) proposed the Adaptive dynamic grouping coding-based (ADG) embedding. They achieved strong statistical concealment. Lu et al. (Lu et al. 2023) used contextual learning and variable length coding-based (VLC) to generate stegos. It improves semantic concealment.
However, the embedding ways of almost all existing schemes allocate one or the same number of codings to the tokens in the candidate pool. That is, when embedding, the probability of selecting all tokens is the same. They do not consider that the CP of a portion of the tokens is not high. If the secret is determined to these tokens, the quality of stego generation will be dropped. This will affect the quality and statistical distribution, and thus the anti-detection ability. The results of the existing embedding ways are given in Figure 1. Therefore, it is necessary to explore a difficult-to-perceive and detect scheme with concealed embedding.
Figure 1: The result of large language model (LLM) + embedding ways. $\uparrow$” and $\downarrow$” represent the higher / lower the value, the better the result. The metrics can be found in “Section 3.1”. The values can be seen in Table 1.
This paper proposes the dynamically allocated interval-based LS scheme, called DAIRstega. Different from the existing LS, DAIRstega introduces the idea of the roulette wheel to design the secret embedding way. We believe that the quality of the tokens in the candidate pool is different. Thus, the tokens need to be given different shoot probabilities by secret, making it more likely that the secret will select tokens with higher CPs. Then, we explore the characteristics of the CP of tokens in the candidate pool, and design a variety of relationships between CPs and the number of allocated codings. By analyzing the adaptability of these allocated functions to the invisible information-hiding tasks of LS, the allocation expression is condensed. In addition, DAIRstega can receive instruction and more directly control the generation of the stego contents. Experiments show that DAIRstega has strong concealment and controllability.
The main contributions are summarized as follows.
  • To augment the concealment of stegos, DAIRstega introduces the idea of the roulette wheel, and dynamically allocates the number of codings. The multiple functions between CPs and allocated numbers are also given. According to Occam’s razor, the functions are condensed.
  • To ensure the adaptability of the functions designed, we gave three constraints that need to be met.
  • To make the evaluation results consistent with the LLMs era, the high-quality corpus was built and used this corpus as a cover to train the model of all schemes.
  • To verify the effectiveness of the proposed embedding way and DAIRstega, we performed experiments regarding various concealment and anti-steganalysis.
  • Experiment verified that DAIRstega can also be regarded as a secure watermarking scheme with high fidelity, consistency, and abuse-resistant.
2 Methodology
2.1 Modeling
For more covert communications, Alice needs to ensure that the discourse (the style of humans and the genre of texts) is the same as the covers, as shown in Figure 2.
Figure 2: Example of texts being transmitted and detected. Responses that match the discourse are difficult to perceive, thus reducing the risk of steganalysis, such as Alice 1.
The covert communication hopes to avoid Eve’s perception and detection and needs to satisfy the following:
$\left\{\begin{array}[]{l}d({P_{C}},{P_{S}})\leq\varepsilon\\ C\Rightarrow\mathcal{C}_{style}\sim{P_{\mathcal{C}1}},{S_{k}}\Rightarrow\mathcal{S}_{style}\sim{P_{\mathcal{S}1}}\\ C\Rightarrow\mathcal{C}_{genre}\sim{P_{\mathcal{C}2}},{\rm{}}{S_{k}}\Rightarrow\mathcal{S}_{genre}\sim{P_{\mathcal{S}2}}\\ (d({P_{\mathcal{C}1}},{P_{\mathcal{S}1}})\leq{\varepsilon_{1}})\wedge(d({P_{\mathcal{C}2}},{P_{\mathcal{S}2}})\leq{\varepsilon_{2}})\end{array}\right.,$ (1)
where, ${P_{C}}$ and ${P_{S}}$ are the distribution of cover and stego, $d(\cdot)$ is the difference, $\varepsilon$ is a very small value greater than 0. $\sim$ is following a certain distribution. The above assumptions cover most situations where the secret is transmitted. Complying with the Formula 1 can improve the concealment and ensure the conduct of covert communication. The results of this aspect are shown in Table 4 and Table 5.
2.2 DAIRstega Overall
This scheme can generate stegos that are strongly concealed and controllable. It can satisfy the extension of covert communication, i.e. Formula 1. To make the generated stegos more consistent with the CP distribution of covers, DAIRstega uses the idea of the roulette wheel to dynamically allocate tokens in the candidate pool to coding intervals of different lengths. That is, the larger the CP is, the larger the roulette area will be allocated, and the greater the shoot that this token will be selected by secret. To make the generated stegos controllable, we introduce the Prompt, taking instructions to control generate. Instruction can be public topics, posted questions, or other requirements in social networks. Figure 3 shows the DAIRstega’s framework.
Figure 3: The DAIRstega’s framework. The “stego generation process” consists of two parts: the “CP generation” and “Embedding” modules. Different from existing LS works, DAIRstega uses the idea of the non-uniform roulette wheel to dynamically allocate different roulette areas (numbers of codings) to tokens in the candidate pool. In this way, the secret is more likely to fall on the tokens with larger CPs, ensuring the quality of stegos. In addition, DAIRstega can not only receive secret information (in the form of bitstreams), but also input the instruction, and finally generate stego that conforms to the instruction.
2.3 Details
Stego generation process.
It is composed of the “CP generation” ($PG$) and “Embedding” ($Em$) modules. The process formulas are as follows:
$\left\{\begin{array}[]{l}PG:LM(I)\Rightarrow\mathbf{P}\\ Em:\text{locate}\;\{{m_{t}}\;\text{in}\;DAI(BC(\mathbf{P}))\}\Rightarrow i,\\ \quad\quad\;\ \text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}\\ i=Em(PG(I))\end{array}\right.,$ (2)
where, $I$ and $\mathbf{P}$ are the index sequence of a given text and the CP distribution of the next token, $\mathbf{P}=\left[{p_{1}},\cdots,{p_{v}}\right]$, ${p_{v}}$ is the CP of the $v$-th token in the vocab. “$LM(\cdot)$” is the CP calculation process. “$DAI(\cdot)$” is the dynamically allocated interval process. “$BC(\cdot)$” is the build candidate pool process. “$\text{locate}\;\{a\;\text{in}\;\circ\}\Rightarrow i$” means that $i$ is the next token determined by $a$. “$\text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}$” is the same prefix that matches the beginning and end of $i$’s range.
CP generation module.
The input is the instruction or the instruction with part of the subsequent token generated, and the output is the CP distribution $\mathbf{P}$ of the next token.
Embedding module.
To ensure that the embedding does not excessively affect the quality of the generated stegos. In our work, low-CP tokens should not be added to the candidate pool. We sort and intercept the first top_k tokens with higher CPs to build the candidate pool. The $\mathbf{P}$ normalization $\mathbf{P}^{\prime}=[{p^{\prime}_{1}},\cdots,{p^{\prime}_{top\_k}}]$ of tokens in candidate pool is:
$p_{j}^{\prime}=\frac{p_{j}}{\sum\nolimits_{j}^{top{\rm{\_}}k}{p_{j}}},p^{\prime}_{j}\in\mathbf{P}^{\prime},p_{j}\in\mathbf{P}.$ (3)
Then, the $\alpha$-bit secret bitstream is read to determine the total number ($12^{\alpha}-$ 1) of intervals. The different numbers of consecutive codings are allocated to tokens in the candidate pool. Intuitively, the allocated relationship between $\mathbf{P}^{\prime}$ and the number of allocated codings $\mathbf{N}=\left[{n_{1}},\cdots,{n_{top\_k}}\right]$ can be linear, the linear function is:
$n_{j}=\lfloor\beta p^{\prime}_{j}\times({2^{\alpha}}-1)\rceil,n_{j}\in\mathbf{N},$ (4)
[…]
Further exploration and discovery, if the linear function is used, the embedding rate can be improved. Tokens with very high $p$ will be allocated to most encodings, and a large interval length will result in a short prefix of this interval regardless of how many bits are read at a time. Longer stego needs to be generated to embed the given bits, resulting in a lower embedding rate (unit: bit per word). So we try to design the nonlinear functions to improve the embedding rate.
Using logarithmic and exponential operations, a variety of nonlinear allocations can be designed. However, not all nonlinear allocations are suitable for LS information hiding. We define three constraints that the functions need to be met:
Constraint 1:
${\frac{{\partial\ {DAI}{(x)}}}{{\partial\ x}}}\geq\frac{c}{1-x},c\in\mathbb{R}^{+},\forall x\in\left[\epsilon^{+},1\right),$ (5)
where, $x$ is a variable that meets all CPs. $\epsilon^{+}$ is a smaller value greater than 0, such as 0.1, 0.2, and so on. Although the CP distribution is discrete, the independent variable of the distribution function is continuous, so it is expressed here using derivatives. This constraint shows that the number of allocations needs to increase with the increase of CP, and ensures that the change rate of the allocated number is not too small. If ${\textstyle{{\partial\;DAI(x)}\over{\partial\;x}}}\leq\varepsilon$, which means the number of allocation tokens is similar. It will degenerate into a similar allocation of the same number of codings as the existing work, thus increasing the selection of a poor token.
Constraint 2:
${\frac{{\partial^{2}{DAI}{(x)}}}{{\partial x^{2}}}}\leq 0,\forall x\in\left[\epsilon^{+},1\right).$ (6)
When the condition is equal, the corresponding function will be transformed into a linear function. This constraint indicates that the entire allocation function should be an increasing function with decreasing growth. If the function has an increasing growth trend, then tokens with larger CPs will be allocated to larger intervals, and the same prefixes at the beginning and end of the interval will be shortened. Thus, the embedding rate will be decreased.
Constraint 3:
$DAI(x)\geq x\;\wedge\;{\frac{{\partial\ {DAI}{(x)}}}{{\partial\ x}}}-{\frac{{\partial\ x}}{{\partial\ x}}}\leq 0,\forall x\in\left[\epsilon^{+},1\right).$ (7)
This constraint shows that the number of codings allocated to tokens with smaller CP should be appropriately increased to increase their probability of being selected. This can also suppress the problem of too large intervals of larger CP to a certain extent.
Here, we give several nonlinear allocation functions:
[…]
According to the principle of Occam’s razor, if you can simplify, you can do less complicated things. we condense the linear and nonlinear functions into a “simple expression” of dynamically allocated intervals that satisfy the above three constraints, as shown below:
$n_{j}=\lfloor(\frac{p^{\prime}}{\sum{p^{\prime}}})^{\beta}\times({2^{\alpha}}-1)\rceil,\beta\in(0,1),\alpha\in\mathbb{Z}^{+}.$ (9)
The smaller $\beta$ is, the smaller the interval of tokens with larger CP will be, while still satisfying the premise that the interval of tokens with larger CP is larger. It will increase the length of the prefix, while reducing the length of stegos to be generated, improving the embedding rate.
Please note that these functions are not fixed, here are just a few examples that meet the constraints. The constraints we proposed are not the most perfect, and the final condensed expression is also for simplicity and easy to understand, and is not the best. This can just provide thinking for subsequent research that is more conducive to hiding information.
Loop the stego generation process, the stego will be obtained. Algorithm 1 illustrates the embedding process.
Algorithm 1 Secret embedding of DAIRstega.
0:  $B=\{x_{0},x_{1},\cdots,x_{m}\},x\in\{0,1\}$; Instruction.
0:  Stego $S=[{t_{n+1}},{t_{n+2}},\cdots,{\rm{<EOS>}}]$.
1:  while Not the end of the bitstream do
2:     Map Instruction $\Rightarrow{I_{1}}=[{i_{1}},{i_{2}},\cdots,{i_{n}}]$;
3:     while Not $I_{<\text{EOS}>}$ in $I$ do
4:        Get $\mathbf{P}$ of $i_{k},k\in[n+1,n+2,\cdots,{\rm{<EOS>}}]$;
5:        Build candidate pool and get $\mathbf{P^{\prime}}$ by Formula 3;
6:        Read the unembedded $\alpha$-bit bitstream;
7:        Give $\beta\Rightarrow n_{j}$ by Formula 9;
8:        Allocate $\mathbf{N}$ to $I$;
9:        Determine the interval by the $\alpha$ bits;
10:     end while
11:     $\text{Locate}\;\{{\alpha\text{-bit}\ B}\;\text{in}\;\text{interval}\}\Rightarrow i_{k}$;
12:     $\text{Match}\;\{i_{k}(R_{begin})\;\text{and}\;i_{k}(R_{end})\}\Rightarrow$ embedded secret at this moment;
13:  end while
14:  Map $[{i_{n+1}},\cdots,{i_{{\rm{<EOS>}}}}]\Rightarrow[{t_{n+1}},\cdots,{\rm{<EOS>}}]$.
Secret extraction process.
Bitstream extraction and embedding are a pair of inverse operations. Same as Discop (Ding et al. 2023b), we also need to use the same scheme, initial context (instruction), and stegos. The instructions remain requirements visible to everyone in social networks. The formula of the extraction process $Ex$ is as follows:
$\left\{\begin{array}[]{l}Ex:\text{locate}\;\ \{{i_{n+1}}\;\ {\rm{in}}\;\ DAI(BC(P))\},\\ \quad\quad\;\text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}\\ B_{j}=Ex(PG(I)),B_{j}\in B\end{array}\right..$ (10)
The interval is located by the stego’s next token at this moment, and the same prefix is the current extracted secret.
3 Experiments
To ensure fairness and reliability in comparisons between schemes, each result was repeated 5 times and averaged to provide the results. All experiments are run on the Linux Server with 24G NVIDIA GeForce RTX 4090 GPUs.
3.1 Settings
Dataset.
We obtained a large amount of data from Wikipedia, Twitter, GPT4 (OpenAI 2023), and the publicly available data for training LLMs111 https://github.com/tloen/alpaca-lora/tree/main. The final dataset contains 108,771 texts and nearly 100 different discourse characteristics to train the language models of all schemes.
Model configuration.
DAIRstega can access any open-source LLMs. In the experiment, we used LLaMA2-7B and 13B models (Meta 2023) as an example. The rank of LoRA (Hu et al. 2021) is 8, and the dropout is 0.05. The learning algorithm is AdamW (Loshchilov and Hutter 2018), and the learning rate is 1e-5. The batch size is 16. For simplicity, we use $\beta=1,0.5$ and $\alpha=8,32,48$ to perform DAIRstega.
Table 1: The overall results of different embedding ways. The embedding rate of each way is near 1Bpw (bit per word). Bold and “ *” are the best and the suboptimal result. “LLM” in this table and Ours are all selected the LLaMA2-7B. For the LS works involved in these embedding ways, see “Section 3.1”. $\uparrow$” is that the higher the value, the better the result. $\downarrow$” means the lower the value, the better the result. The complete data can be seen in Table 8 and Table 9 in Supplementary A.2.
[…]
PPL $\downarrow$ $\Delta$Pcs $\downarrow$ CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$ MD $\downarrow$ $\Delta$DP $\downarrow$ Mauve $\uparrow$ BLEU $\uparrow$
LLM+HC 13.331 5.639 96.54 45.79 0.263 8.255* 3.46 70.60 54.26
LLM+ADG 12.750* 5.058* 97.20* 45.59* 0.240* 8.316 2.80* 77.86 71.79*
LLM+AC 73.454 65.762 82.16 51.64 0.746 15.415 17.84 35.64 50.92
LLM+FLC 12.883 5.190 97.04 46.56 0.241 8.280 2.96 79.48 54.88
Ours 8.030 0.342 97.43 38.60 0.227 6.341 2.57 78.53* 77.42
Embedding $\hookrightarrow$ Anti-steganalysis
Score $\uparrow$ LS_CNN TS_CSW EILG UP4LS
Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$
LLM+HC 58.44 77.62 78.41 78.09 78.25 78.58 78.29 86.73 86.13
LLM+ADG 59.18* 70.10* 70.17* 69.45* 69.34* 70.99* 70.49* 84.00* 84.23*
LLM+AC 45.10 90.15 89.98 89.69 89.37 86.92 86.52 93.62 93.37
LLM+FLC 58.44 78.53 80.05 82.72 83.32 76.67 76.15 89.03 89.20
Ours 65.41 58.95 58.74 55.28 61.69 65.01 65.83 72.92 73.30
Baselines.
Comparison is divided into two parts, one is the comparison between the proposed embedding way and the existing embedding ways, and the other is the comparison between DAIRstega and the prevalent works in recent years.
$\cdot$ In terms of the embedding ways, the existing LS works can be summarized as follows: (1) HC (or VLC), (2) ADG, (3) AC, and (4) FLC. (1) HC is used in VAEstega (Yang et al. 2021), LSCS (Lu et al. 2023), Hi-Stega (Wang et al. 2023a), RNNstega (Yang et al. 2019), GAN-APD (Zhou et al. 2021), JLS (Ding et al. 2023a), Discop (Ding et al. 2023b). (2) ADG is used in Hi-Stega (Wang et al. 2023a), ADG (Zhang et al. 2021). (3) AC is used in VAEstega (Yang et al. 2021), Hi-Stega (Wang et al. 2023a). (4) FLC is used in RNNstega (Yang et al. 2019), JLS (Ding et al. 2023a).
$\cdot$ In terms of the prevalent works, we selected the LS and watermarking works that showed excellent performance as the baselines. LS: (1) Fang (ACL) (Fang, Jaggi, and Argyraki 2017). (2) RNNstega (IEEE TIFS) (Yang et al. 2019). (3) ADG (ACL-IJCNLP) (Zhang et al. 2021). (4) LSCS (IJCNN) (Lu et al. 2023). Watermarking: (5) PLMmark (AAAI) (Liu et al. 2023). It uses pre-trained language models to provide the first secure black-box watermarking. They can compare the generation effects of DAIRstega from different aspects, and these performances have been widely recognized.
Evaluation metrics.
$\downarrow$ means the lower the value, the better, $\uparrow$ means the higher the value, the better.
$\cdot$ In terms of text quality (perceptual concealment), we adopt (1) PPL $\downarrow$ (stego perplexity). (Wu et al. 2023; Ding et al. 2023a). (2) $\Delta$Pcs $\downarrow$ (The difference between PPL of cover and stego (Yang et al. 2021)).
$\cdot$ In terms of statistical analysis (statistical concealment), we adopt (3) CS $\uparrow$. (Cosine similarity). (4) JSD $\downarrow$, which averages the KL divergence. (5) ED $\downarrow$ (Euclidean distance). (6) MD $\downarrow$ (Manhattan distance). (7) $\Delta$DP $\downarrow$ (Dot product difference (Cai et al. 2021)).
$\cdot$ In terms of discourse matching (Semantic concealment), we adopt (8) $\Delta$LDA $\downarrow$. (9) Mauve $\uparrow$ (Pillutla et al. 2021). Mauve is a metric that considers fluency and consistency. (10) BLEU $\uparrow$ (Bilingual evaluation understudy (Papineni et al. 2002)). (11) Score (BERTScore) $\uparrow$ (Zhang et al. 2020). Score considers the similarity at the segment level.
$\cdot$ In terms of anti-steganalysis, we use the high-performance linguistic steganalysis LS_CNN (Wen et al. 2019), TS_CSW (Yang, Huang, and Zhang 2020), EILG (Xu, Zhang, and Liu 2023), and UP4LS (Wang et al. 2023b): (12) Acc $\downarrow$ (Accuracy) and (13) F1 $\downarrow$ (F1-score).
3.2 Comparison of the embedding ways
To illustrate the effect of embedding way, we performed this experiment. We add the existing works with the same LLMs as ours. The results are shown in Table 1.
The results in Table 1 show that the proposed embedding way can improve the stego quality, better simulate the CP distribution of covers, and get the optimal anti-steganalysis. Overall performance performs better than the existing ways.
3.3 Comparison with LS baselines
Text quality (Perceptual concealment).
Table 2 shows the comparison regarding the stego quality.
Table 2: Text quality comparison of stegos generated by DAIRstega and LS baselines. Bold is the best result. “ *” is the suboptimal result. $\alpha$ and $\beta$ can be found in Formula 9. The meanings of the parameters “bin”, “bit”, and “$\tau$” are the same in Fang, RNNstega, and ADG schemes.
Schemes Param / Bpw PPL $\downarrow$ $\Delta$Pcs $\downarrow$
Fang bin = 1 / 1.00 235.271 227.578
bin = 3 / 3.00 269.303 261.611
RNNstega bit = 1 / 1.00 79.987 72.294
bit = 3 / 2.63 99.302 91.610
ADG $\tau=0.5$ / 4.38 197.498 189.806
LSCS - / 1.12 9.248 1.556
DAIRstega $\beta$ = 1 $\alpha$ = 8 / 1.10 7.686 0.006
$\alpha$ = 32 / 1.11 8.184* 0.492*
$\alpha$ = 48 / 1.13 8.219 0.527
[…]
$\alpha$ = 32 / 2.56 20.876 13.184
$\alpha$ = 48 / 2.58 22.069 14.377
According to the results in Table 2, it can be found that the stego quality of DAIRstega markedly surpasses that of the baselines. This shows that DAIRstega can be more consistent with the linguistic characteristics of covers, enhancing the perceptual concealment of the stegos.
Statistical analysis (Statistical concealment).
Table 3 shows the comparison between DAIRstega and baselines regarding the statistical analysis between covers and stegos.
Table 3: Statistical analysis comparison of stegos generated by DAIRstega and LS baselines. The meanings of “Bold”, “ *”, “bin”, “bit”, and “$\tau$” are the same in Table 2.
Schemes Param CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$ MD $\downarrow$ $\Delta$DP $\downarrow$
Fang bin = 1 9.31 63.43 1.347 17.911 90.69
bin = 3 9.30 63.34 1.347 17.876 90.70
RNNstega bit = 1 95.97 41.51 0.284 7.530 4.03
bit = 3 96.32 39.38 0.271 7.371 3.68
ADG $\tau=0.5$ 96.82 43.02 0.252 8.232 3.18
LSCS - 94.20 40.28 0.340 7.182 5.80
DAIRstega $\beta$=1 $\alpha$=8 97.20 38.86 0.237 6.458 2.80
$\alpha$=32 97.65 38.48 0.217 6.293* 2.35
$\alpha$=48 97.45* 38.46 0.226* 6.273 2.55*
[…]
$\alpha$=32 92.43 39.69 0.389 8.976 7.57
$\alpha$=48 92.44 38.31* 0.389 8.592 7.56
According to the comparison results in Table 3, it can be found that DAIRstega can better simulate the statistical distribution of cover, enhancing the statistical concealment.
Discourse matching (Semantic concealment).
Table 4 shows the comparison regarding discourse matching.
Table 4: Discourse-matching comparison of stegos generated by DAIRstega and LS baselines. Unit is %. The complete data can be seen as Table 11 in Supplementary A.3.
Schemes Param $\Delta$LDA $\downarrow$ Mauve $\uparrow$ BLEU $\uparrow$ Score $\uparrow$
Fang bin = 1 11.964 2.15 0.90 42.58
bin = 3 11.964 1.96 0.90 41.65
RNNstega bit = 1 0.062 12.73 1.87 52.57
bit = 3 18.919 11.97 1.83 52.36
ADG $\tau=0.5$ 15.901 18.54 2.72 47.55
LSCS - 0.022 2.40 31.95 59.22
DAIRstega $\beta$=1 $\alpha$=8 0.017* 79.91* 76.37 65.76
$\alpha$=32 0.022 76.97 77.85* 65.22
$\alpha$=48 0.020 78.70 76.53 65.25*
[…]
$\alpha$=32 0.029 75.58 66.99 60.89
$\alpha$=48 0.006 72.40 79.64 60.29
Anti-steganalysis
Table 5 shows the anti-steganalysis comparison of DAIRstega and baselines.
Table 5: Anti-steganalysis (Acc $\downarrow$) comparison of stegos generated by DAIRstega and LS baselines. Unit is %. The complete data can be seen as Table 12 in Supplementary A.3. We give the detection of the other 7 steganalysis (including non-BERT, BERT, and LLMs-based). It can be seen as Table A.4 in Supplementary A.4.
Schemes Param LS_CNN TS_CSW EILG UP4LS
Fang bin = 1 99.70 88.34 99.50 99.75
bin = 3 99.75 89.03 99.52 99.88
RNNstega bit = 1 85.81 81.39 85.86 97.42
bit = 3 82.93 80.35 82.71 96.78
ADG $\tau=0.5$ 85.86 82.23 84.86 98.34
LSCS - 78.86 81.19 84.15 96.22
DAIRstega $\beta$=1 $\alpha$=8 59.50 56.82 67.25 75.40
$\alpha$=32 59.35* 54.05 62.28 71.97
$\alpha$=48 58.00 54.98* 65.51* 71.40*
[…]
$\alpha$=32 72.06 68.73 71.71 76.41
$\alpha$=48 70.17 69.48 70.64 79.46
3.4 Comparison with watermarking baseline
The rise of LLMs has spawned many excellent watermarking papers (Liu et al. 2023; Kirchenbauer et al. 2023). They provide solutions for copyright protection in terms of high fidelity (Kirchenbauer et al. 2023), consistency (Zhang et al. 2024), and abuse resistance (Liu et al. 2023). Excitingly, DAIRstega can also be regarded as an invisible watermarking with strong fidelity, consistency, and abuse resistance.
The focus of this paper is not watermarking, we only compare it with the PLMmark (AAAI 2023), as shown in Table 6.
Table 6: Fidelity, consistency, and abuse resistance comparison with watermarking baseline. The complete data can be seen in Table 11 and Table 12 in Supplementary A.3.
[…]
PPL $\downarrow$ $\Delta$Pcs $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$
PLMmark 18.748 11.056 0.009 7.22
DAIRstega 7.686 0.006 0.006 80.83
Scheme $\hookrightarrow$ Abuse resistance
Score $\uparrow$ TS_CSW $\downarrow$ EILG $\downarrow$ UP4LS $\downarrow$
PLMmark 43.50 98.81 99.92 99.85
DAIRstega 65.76 54.05 62.28 69.04
According to the comparison results in Table 6, the watermark texts generated by DAIR have excellent fidelity and consistency, and it is difficult for unauthorized receivers to detect the existence of the watermark text, avoiding the problem of watermark abuse. Only authorized receivers with the same model settings can detect and fully extract the watermark information.
3.5 Longer stegos by DAIRstega
Table 7: Examples of long stegos generated by DAIRstega. Bitstream is the binary of the secret. Secret can be not only texts, but also other modalities. In addition, we embed the “AAAI 2025 icon”. For details, see Table 14 in Supplementary A.5.
Secret&Bitsream Instruction Stegos PPL $\Delta$LDA
Secret:
Love and peace
Bitstream:
010011000110 111101110110 011001010010 000001100’00 101101110011 001000010000 001110000011 00101’011000 010110001101 100101 (114bit)
Discuss the causes of the Great Depression.
The Great Depression was a severe economic downturn that lasted from 1929 to 1939. It was the worst economic depression in the history of the United States and it affected many countries around the world. The causes of the Great Depression are numerous and complex, but some of the main factors include overproduction, the bursting of the stock market bubble, banking failures, and government policies. One of the primary causes of the Great Depression was overproduction. The United States economy was booming in the 1920s, and many businesses and industries expanded their production to meet the growing demand. However, this resulted in an oversupply of goods and services, which led to a decrease in prices and a decrease in demand. This decrease in demand caused a drop in profits, which led to businesses cutting back on production and laying off workers.
2.145
0.004
Explain the use of word embeddings in Natural Language Processing. Word embeddings are a form of representation learning in natural language processing (NLP). Word embeddings are vectors of real numbers that represent the meaning of words. These vectors are learned from large amounts of text data and can be used to model the relationship between words and their contexts. Word embeddings can be used in various NLP tasks such as sentiment analysis, machine translation, and text classification.
3.399
0.006
Give some strategies for getting over a fear of public speaking. Public speaking is a common fear for many people, and it can be difficult to overcome. Here are some strategies that can help: 1. Practice: The more you practice public speaking, the more comfortable you will become. Start by practicing in front of a small group of friends or family members, and gradually work your way up to larger audiences. 2. Visualize: Before giving a speech, visualize yourself giving a successful presentation. Imagine yourself confidently delivering your speech, and see yourself receiving positive feedback from the audience. 3. Breathe: Take deep, slow breaths before giving a speech.
2.046
0.007
The secret will change the next token. This makes it difficult to maintain logical coherence when generating longer stegos. So longer stegos with high quality is a problem, and there is little research in this task. Fortunately, we found that DAIRstega also has this ability. Examples of embedded text and cross-modal secrets are shown in Table 7 and Table 14.
3.6 Ablation experiment
We performed ablation experiment on different LLMs, and the results are shown in Table 8.
Table 8: Ablation on different LLMs. Bold is the best result. Here we only take “$\beta$ = 1, $\alpha$ = 32” as an example.
$\beta$=1, $\alpha$=32 PPL $\downarrow$ CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$
LLaMA2-7B 8.184 97.65 38.48 0.217
LLaMA2-13B 8.431 96.32 43.93 0.271
$\beta$=1, $\alpha$=32 MD $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$ BLEU $\uparrow$
LLaMA2-7B 6.293 0.022 76.97 77.85
LLaMA2-13B 8.034 0.019 88.38 72.31
$\beta$=1, $\alpha$=32 $\Delta$DP $\downarrow$ Score $\uparrow$ TS_CSW $\downarrow$ UP4LS $\downarrow$
LLaMA2-7B 2.35 65.22 54.05 71.97
LLaMA2-13B 3.68 61.66 54.73 71.12
According to Table 8, it is not possible to generate stegos with stronger concealment by using larger LLMs. This is because LS needs to embed secret. The stego quality is more relevant with embedding way rather than using larger LLMs.
4 Conclusion
To enhance the concealment of stegos, this paper proposes an LS scheme based on dynamically allocated intervals, namely DAIRstega. It uses the idea of non-uniform roulette and takes the CP of tokens as an important basis for allocating the number of codings. Then the requirements of invisible information hiding are analyzed and three constraints are given. The allocation expression is condensed. Experiments show that the proposed embedding way and DAIRstega have achieved excellent performance in various concealment and anti-steganalysis. We also prove that the stego quality depends more on the design of the model and embedding way, not just the use of LLMs. Furthermore, DAIRstega can be used as a secure watermarking scheme. Last but not least, to concisely show the proposed scheme, we did not conduct the results of all functions and LLMs.
5 Outlook
We have stimulated thinking in this field and discovered gaps in the field for scholars to conduct follow-up research.
1. Given the limited studies on longer and more diverse LS. The scheme can be designed to improve the length and diversity. 2. We know that fine-tuning LLMs brings huge costs, so the fine-tuning-free LS can be improved. Knowledge Graphs or Chain of Thought are used to enhance the prompts, and encoding text features to achieve a convenient and completely concealed LS scheme. 3. For watermarking, the idea of LS design can be used to take the hyperparameters of the open-source LLMs as the part that needs to be verified, and construct a new watermarking scheme. 4. Virtual social robots can be built, leveraging secrets to generate multi-turn conversations. It promises contactless covert communication, keeping the sender safe. 5. We recognize the shortcomings of current steganalysis for detecting DAIRstega’s stegos. So steganalysis needs to be developed.
Appendix A Supplementary
A.1 Related work
The concealment of stego determines the success of covert communication to a great extent. The concealment is primarily manifested in three aspects: perceptual, statistical, and semantic.
Perceptual concealment” focuses on ensuring that the scheme generates stego with complete and fluent. This is the most basic requirement of LS. In pursuit of this objective, Fang et al. (Fang, Jaggi, and Argyraki 2017) proposed an LSTM-based LS scheme. This scheme segments the vocabulary into several sets based on bit blocks. It then selects tokens with the highest probability that corresponds to the secret information from the candidate pool. Ding et al. (Ding et al. 2023a) combined the conditional generation strategy with the replacement technique, using text sequences as auxiliary data in the stego generation process to enhance the embedding capabilities. Yang et al. (Yang et al. 2019) trained a language model to generate stegos. Using HC-based (Huffman coding) and FLC-based (Fixed length coding) to encode the candidate pool, these stegos have excellent completeness and fluentness.
Statistical concealment” requires the stegos distribution to be close to the covers. To achieve this goal, Yang et al. (Yang et al. 2021) designed an encoder-decoder structure, and VLC-based (Variable length coding) is used to encode the CP. To further reduce the distribution difference between cover and stego, Zhang et al. (Zhang et al. 2021) used adaptive dynamic grouping coding to recursively embed secret. Zhou et al. (Zhou et al. 2021) used a GAN (Generative Adversarial Network) to design an adaptive probability distribution steganography. These ensure statistical concealment.
Semantic concealment” aims to generate the stego that is coherent and semantically controllable. To this end, Li et al. (Li et al. 2021) put forward an LS scheme based on the knowledge graph. Yang et al. (Yang et al. 2023) utilized semantic information encoding to embed secret information, realizing the effect of maintaining semantics and increasing the embedding capacity during the translation process. Wang et al. (Wang et al. 2023a) leveraged the relevance of social network context to enhance contextual semantic relevance while maintaining embedding rates. Lu et al. (Lu et al. 2023) used contextual learning and GPT2 guidance to generate stegos.
A.2 Supplement the effect of the proposed embedding way
Table 9: Text quality, statistical analysis, and discourse-matching comparison of stegos generated by DAIRstega and baselines. The embedding rates of each way include 1 Bpw (bit per word) and 2 Bpw. “Avg.” represents the overall performance. Bold is the best overall result. “ *” is the suboptimal overall result. The meanings of “bit” are the same in Table 2.
Embedding Param / Bpw Text quality Statistical analysis Discourse-matching
PPL $\downarrow$ $\Delta$Pcs $\downarrow$ CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$ MD $\downarrow$ $\Delta$DP $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$ BLEU $\uparrow$ Score $\uparrow$
LLMs+HC bit = 1 / 1.00 13.331 5.639 96.54 45.79 0.263 8.255 3.46 0.014 70.60$\pm$27.36 54.26$\pm$4.25 58.44$\pm$4.07
bit = 2 / 2.52 24.334 16.641 90.57 47.35 0.350 9.354 9.43 0.009 64.88$\pm$30.03 55.00$\pm$5.77 56.19$\pm$4.01
Avg. 18.832 11.140 93.55 46.57 0.307 8.805 6.45 0.011 67.74 54.63 57.32
LLMs+ADG $\tau$ = 0.3 / 1.34 12.750 5.058 97.20 45.59 0.240 8.316 2.80 0.010 77.86$\pm$23.65 71.79$\pm$3.03 59.18$\pm$5.84
$\tau$ = 0.65 / 2.99 16.175 8.483 95.95 44.74 0.331 8.478 4.05 0.018 75.60$\pm$27.02 62.60$\pm$4.68 58.60$\pm$5.69
Avg. 14.463* 6.770* 96.58 45.17* 0.285 8.397* 3.42 0.014 76.73* 67.20* 58.35*
LLMs+AC CP = 8 / 1.25 73.454 65.762 82.16 51.64 0.746 15.415 17.84 0.014 35.64$\pm$22.65 50.92$\pm$29.30 45.10$\pm$3.07
CP = 32 / 1.32 86.238 78.545 80.38 51.21 0.829 14.194 19.62 0.012 37.18$\pm$24.59 48.01$\pm$27.45 47.51$\pm$2.76
Avg. 79.846 72.154 81.27 51.43 0.788 14.805 18.73 0.013* 36.41 49.47 46.31
LLMs+FLC bit = 1 / 1.34 12.883 5.190 97.04 46.56 0.241 8.280 2.96 0.031 79.48$\pm$26.53 54.88$\pm$6.33 58.44$\pm$4.64
bit = 2 / 2.67 30.481 22.788 93.86 46.30 0.328 8.939 6.14 0.011 61.10$\pm$29.30 58.21$\pm$9.88 55.20$\pm$3.39
Avg. 21.682 13.989 95.45 46.43 0.284* 8.610 4.55 0.021 70.29 56.55 56.82
Ours $\beta$=1 $\alpha$=8 / 1.10 7.686 0.006 97.20 38.86 0.237 6.458 2.80 0.017 79.91$\pm$26.75 77.88$\pm$30.40 65.76$\pm$4.13
$\beta$=1 $\alpha$=32 / 1.11 8.184 0.492 97.65 38.48 0.217 6.293 2.35 0.022 76.97$\pm$26.47 77.85$\pm$30.42 65.22$\pm$4.61
$\beta$=1 $\alpha$=48 / 1.13 8.219 0.527 97.45 38.46 0.226 6.273 2.55 0.020 78.70$\pm$25.51 76.53$\pm$31.71 65.25$\pm$4.78
Avg. (1Bpw) 8.030 0.342 97.43 38.60 0.227 6.341 2.57 0.020 78.53 77.42 65.41
[…]
$\beta$=0.5 $\alpha$=32 / 2.56 20.876 13.184 92.43 39.69 0.389 8.976 7.57 0.029 75.58$\pm$20.11 66.99$\pm$27.08 60.89$\pm$3.85
$\beta$=0.5 $\alpha$=48 / 2.58 22.069 14.377 92.44 38.31 0.389 8.592 7.56 0.006 72.40$\pm$26.33 79.64$\pm$12.47 60.29$\pm$4.00
Avg. (2Bpw) 18.907 11.214 94.09 38.60 0.336 8.258 5.91 0.018 76.27 73.27 61.22
[…]
Table 10: Anti-steganalysis comparison of stegos generated by DAIRstega and baselines. The embedding rates of each way include 1Bpw (bit per word) and 2Bpw. “Avg.” represents the overall performance. Bold is the best overall result. “ *” is the suboptimal overall result.
Embedding Param / Bpw LS_CNN TS_CSW EILG UP4LS
Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$
LLMs+HC bit = 1 / 1.00 77.62$\pm$1.38 78.41$\pm$1.65 78.09$\pm$1.33 78.25$\pm$1.21 78.58$\pm$1.69 78.29$\pm$2.24 86.73$\pm$1.32 86.13$\pm$0.50
bit = 2 / 2.52 77.54$\pm$1.50 77.57$\pm$2.05 76.03$\pm$1.74 76.09$\pm$1.33 78.29$\pm$0.72 78.95$\pm$1.47 88.19$\pm$1.00 88.43$\pm$2.61
Avg. 77.58 77.99 77.06 77.17 78.44 78.62 87.46 87.28
LLMs+ADG $\tau$ = 0.3 / 1.34 70.10$\pm$0.66 70.17$\pm$0.89 69.45$\pm$2.05 69.34$\pm$2.53 70.99$\pm$3.15 70.49$\pm$1.11 84.00$\pm$0.83 84.23$\pm$1.07
$\tau$ = 0.65 / 2.99 63.06$\pm$2.49 63.69$\pm$3.61 62.22$\pm$1.46 61.86$\pm$1.18 66.14$\pm$1.38 66.15$\pm$1.38 80.42$\pm$0.59 80.57$\pm$0.58
Avg. 66.58* 66.93* 65.84* 65.60* 68.57* 68.32* 82.21* 82.40*
LLMs+AC CP = 8 / 1.25 90.15$\pm$1.17 89.98$\pm$1.21 89.69$\pm$0.93 89.37$\pm$0.88 86.92$\pm$1.12 86.52$\pm$0.96 93.62$\pm$0.89 93.37$\pm$0.40
CP = 32 / 1.32 88.45$\pm$1.38 88.44$\pm$1.25 90.80$\pm$1.97 90.52$\pm$2.18 86.44$\pm$2.47 85.86$\pm$3.10 94.03$\pm$0.64 94.16$\pm$0.95
Avg. 89.30 89.21 90.25 89.95 86.68 86.19 93.83 93.77
LLMs+FLC bit = 1 / 1.34 78.53$\pm$1.20 80.05$\pm$1.39 82.72$\pm$1.27 83.32$\pm$1.10 76.67$\pm$0.85 76.15$\pm$1.96 89.03$\pm$0.25 89.20$\pm$0.80
bit = 2 / 2.67 76.49$\pm$1.15 76.54$\pm$1.13 77.08$\pm$2.06 77.26$\pm$1.75 76.95$\pm$1.69 76.46$\pm$1.74 90.62$\pm$0.89 90.37$\pm$0.40
Avg. 77.51 78.30 79.90 80.29 76.81 76.31 89.83 89.79
Ours $\beta$=1 $\alpha$=8 / 1.10 59.50$\pm$2.77 60.15$\pm$10.59 56.82$\pm$4.09 64.21$\pm$10.03 67.25$\pm$3.23 68.04$\pm$1.53 75.40$\pm$0.37 76.26$\pm$1.28
$\beta$=1 $\alpha$=32 / 1.11 59.35$\pm$1.25 59.99$\pm$5.51 54.05$\pm$4.72 61.68$\pm$11.17 62.28$\pm$0.50 62.66$\pm$2.31 71.97$\pm$0.64 72.75$\pm$1.21
$\beta$=1 $\alpha$=48 / 1.13 58.00$\pm$2.37 56.08$\pm$13.79 54.98$\pm$3.14 59.17$\pm$11.37 65.51$\pm$2.23 66.78$\pm$3.91 71.40$\pm$0.65 70.89$\pm$2.05
Avg. (1Bpw) 58.95 58.74 55.28 61.69 65.01 65.83 72.92 73.30
[…]
$\beta$=0.5 $\alpha$=32 / 2.56 72.06$\pm$1.17 71.78$\pm$1.21 68.73$\pm$1.29 68.19$\pm$3.44 71.71$\pm$1.95 70.34$\pm$2.41 76.41$\pm$2.56 75.90$\pm$2.73
$\beta$=0.5 $\alpha$=48 / 2.58 70.17$\pm$1.81 70.40$\pm$3.11 69.48$\pm$0.68 67.80$\pm$1.98 70.64$\pm$1.61 71.00$\pm$2.23 79.46$\pm$0.56 78.28$\pm$1.13
Avg. (2Bpw) 68.37 68.13 67.11 65.25 69.34 69.06 74.97 74.14
[…]
A.3 Complete data for Table 4 and Table 5
In this section, we give the complete data of Table 4 and Table 5, as shown in Table 11 and Table 12. Table 11 contains the standard deviations of many datasets with different discourse, and Table 12 contains the standard deviations of multiple runs.
Table 11: Complete data for Table 4. The meanings of “bin” and “bit” are the same in Table 2. “a±$\pm$±b” represents “Avg.$\pm$±Std”. Bold is the best result. “ *” is the suboptimal result. The “Std”s of DAIRstega Mauve and BLEU values are bigger. This is not due to the instability of 5-time runs, but is obtained from the effect on dozens of different discourse datasets, and represents the overall performance.
Schemes Param / Bpw Mauve $\uparrow$ BLEU $\uparrow$ Score $\uparrow$
Fang bin = 1 / 1.00 2.15$\pm$0.87 0.90$\pm$1.58 42.58$\pm$1.30
bin = 3 / 3.00 1.96$\pm$0.68 0.90$\pm$1.59 41.65$\pm$1.33
RNNstega bit = 1 / 1.00 12.73$\pm$10.56 1.87$\pm$3.28 52.57$\pm$2.20
bit = 3 / 2.63 11.97$\pm$10.56 1.83$\pm$3.21 52.36$\pm$1.80
ADG $\tau$ = 0.5 / 4.38 18.54$\pm$13.89 2.72$\pm$4.79 47.55$\pm$2.25
LSCS - / 1.12 2.40$\pm$1.78 31.95$\pm$12.89 59.22$\pm$3.14
PLMmark - / - 7.22$\pm$4.88 0.15$\pm$0.27 43.50$\pm$0.79
DAIRstega $\beta$ = 1 $\alpha$ = 8 / 1.10 79.91$\pm$26.75 76.37$\pm$31.85 65.76$\pm$4.13
$\alpha$ = 32 / 1.11 76.97$\pm$26.47 77.85$\pm$30.42* 65.22$\pm$4.61
$\alpha$ = 48 / 1.13 78.70$\pm$25.51 76.53$\pm$31.71 65.25$\pm$4.78*
[…]
$\alpha$ = 32 / 2.56 75.58$\pm$20.11 66.99$\pm$27.08 60.89$\pm$3.85
$\alpha$ = 48 / 2.58 72.40$\pm$26.33 79.64$\pm$12.47 60.29$\pm$4.00
Table 12: Complete data for Table 5. The meanings of “bin” and “bit”are the same in Table 2. “a±$\pm$±b” represents “Avg.$\pm$±Std”. “Avg” and “Std” in this table are obtained by running 5 times. Bold is the best result. “ *” is the suboptimal result.
Schemes Param / Bpw LS_CNN TS_CSW EILG UP4LS
Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$
Fang bin = 1 / 1.00 99.70$\pm$0.29 99.68$\pm$0.31 88.34$\pm$21.22 91.01$\pm$15.69 99.50$\pm$0.25 99.51$\pm$0.25 99.75$\pm$0.06 99.60$\pm$0.09
bin = 3 / 3.00 99.75$\pm$0.16 99.76$\pm$0.15 89.03$\pm$20.58 92.39$\pm$13.82 99.59$\pm$0.17 99.52$\pm$0.66 99.88$\pm$0.10 99.80$\pm$0.15
RNNstega bit = 1 / 1.00 85.81$\pm$1.93 86.03$\pm$1.40 81.39$\pm$1.70 82.59$\pm$1.62 85.86$\pm$0.50 86.29$\pm$0.89 97.42$\pm$0.44 95.97$\pm$0.67
bit = 3 / 2.63 82.93$\pm$2.18 84.48$\pm$2.20 80.35$\pm$2.38 77.94$\pm$3.04 82.71$\pm$0.17 83.10$\pm$1.02 96.78$\pm$0.82 94.74$\pm$1.34
ADG $\tau$ = 0.5 / 4.38 85.86$\pm$1.73 86.35$\pm$1.82 82.23$\pm$2.40 82.08$\pm$2.96 84.86$\pm$1.74 84.10$\pm$1.80 98.34$\pm$0.39 97.41$\pm$0.58
LSCS - / 1.12 78.86$\pm$3.07 78.98$\pm$2.83 81.19$\pm$6.62 80.11$\pm$8.26 84.15$\pm$0.41 85.11$\pm$0.79 96.22$\pm$2.08 95.48$\pm$3.11
PLMmark - / - 98.64$\pm$0.92 98.60$\pm$0.95 98.81$\pm$1.83 98.75$\pm$1.93 99.92$\pm$0.17 99.92$\pm$0.11 99.85$\pm$0.30 99.85$\pm$0.30
DAIRstega $\beta$ = 1 $\alpha$ = 8 / 1.10 59.50$\pm$2.77 60.15$\pm$10.59 56.82$\pm$4.09 64.21$\pm$10.03 67.25$\pm$3.23 68.04$\pm$1.53 75.40$\pm$0.37 76.26$\pm$1.28
$\alpha$ = 32 / 1.11 59.35$\pm$1.25 59.99$\pm$5.51* 54.05$\pm$4.72 61.68$\pm$11.17 62.28$\pm$0.50 62.66$\pm$2.31* 71.97$\pm$0.64 72.75$\pm$1.21
$\alpha$ = 48 / 1.13 58.00$\pm$2.37 56.08$\pm$13.79 54.98$\pm$3.14* 59.17$\pm$11.37 65.51$\pm$2.23 66.78$\pm$3.91 71.40$\pm$0.65* 70.89$\pm$2.05*
[…]
$\alpha$ = 32 / 2.56 72.06$\pm$1.17 71.78$\pm$1.21 68.73$\pm$1.29 68.19$\pm$3.44 71.71$\pm$1.95 70.34$\pm$2.41 76.41$\pm$2.56 75.90$\pm$2.73
$\alpha$ = 48 / 2.58 70.17$\pm$1.81 70.40$\pm$3.11 69.48$\pm$0.68 67.80$\pm$1.98 70.64$\pm$1.61 71.00$\pm$2.23 79.46$\pm$0.56 78.28$\pm$1.13
A.4 Supplement data for anti-steganalysis
We also perform other 7 steganalysis methods to detect the stegos we generated, as shown in Table A.4. These include non-BERT-based, BERT-based, and LLMs-based steganalysis methods.
Table 13: The supplemental anti-steganalysis capability of the DAIRstega from other 7 steganalysis works. Bold represents the best result and the best Avg. result. “ *” represents the suboptimal result.
Supplemental VAE-Stega [1] DAIRstega
[…]
non-BERT- [2] Acc $\downarrow$ 57.63 53.75 52.55 61.63 57.75 74.38 59.61 54.38 51.37 51.60* 56.97 52.35 59.50 54.38
F1 $\downarrow$ 50.65 34.63 46.31* 60.89 66.65 73.89 55.50 47.09 47.55 49.47 52.35 49.20 57.70 50.56
based [3] Acc $\downarrow$ 64.50 58.75 62.50 87.63 80.75 94.38 74.75 64.75 61.05* 61.88 67.42 63.83 68.73 64.61
F1 $\downarrow$ 66.91 56.46 64.95 88.03 82.14 94.33 75.47 63.85 61.05* 67.03 66.16 63.82 68.83 65.12
BERT- [4] Acc $\downarrow$ 87.38 74.03 92.25 92.75 86.81 97.38 88.43 70.63 65.43* 65.06 65.87 65.56 73.25 67.63
F1 $\downarrow$ 87.39 74.21 91.99 93.21 86.85 97.35 88.50 70.44 61.48 65.28 65.15* 67.53 74.27 67.36
[5] Acc $\downarrow$ 90.75 78.75 95.25 95.25 88.38 98.13 91.08 70.38 65.50* 64.38 71.12 65.50* 74.25 68.52
based F1 $\downarrow$ 90.74 78.75 95.25 95.25 88.36 98.12 91.08 70.34 65.50 64.27 70.99 65.31* 74.25 68.44
[6] Acc $\downarrow$ 92.50 75.38 95.88 94.75 88.32 97.62 90.74 70.12 64.25 64.25 73.75 66.50* 73.88 68.79
F1 $\downarrow$ 92.19 73.20 95.62 94.53 88.26 97.60 90.23 72.31 66.19 66.29* 71.47 70.02 69.04 69.22
LLMs- [7] Acc $\downarrow$ 89.63 78.13 93.13 88.63 76.38 86.75 85.44 63.38 59.63* 61.44 63.25 59.38 62.13 61.53
F1 $\downarrow$ 89.59 78.06 93.12 88.62 76.13 86.75 85.38 63.24 60.15 60.59* 63.13 59.15 61.86 61.35
based [8] Acc $\downarrow$ 97.38 92.38 98.38 98.13 93.88 98.88 96.50 75.38 72.63 71.75* 81.75 69.38 83.25 75.69
F1 $\downarrow$ 96.87 92.97 98.57 97.10 92.97 98.97 96.25 75.27 72.41 71.35* 81.55 69.32 83.15 75.51
  • [1] VAE-Stega: Linguistic Steganography Based on Variational Auto-Encoder. IEEE Trans. Inf. Forensics and Security, 880–895, 2021.
  • [2] A Fast and Efficient Text Steganalysis Method. IEEE Signal Processing Lett., 26(4): 627–631, 2019.
  • [3] A Hybrid R-BiLSTM-C Neural Network Based Text Steganalysis. IEEE Signal Processing Lett., 26(12): 1907–1911, 2019.
  • [4] High-performance Linguistic Steganalysis, Capacity Estimation and Steganographic Positioning. In Proceeding of IWDW, 2021.
  • [5] Small-Scale Linguistic Steganalysis for Multi-Concealed Scenarios. IEEE Signal Processing Lett., 29: 130–134, 2022.
  • [6] SeSy: Linguistic Steganalysis Framework Integrating Semantic and Syntactic Features. IEEE Signal Processing Lett., 31–35, 2021.
  • [7] Towards Next-Generation Steganalysis: LLMs Unleash the Power of Detecting Steganography. arXiv:2405.09090, 2024.
  • [8] Linguistic Steganalysis via LLMs: Two Modes for Efficient Detection of Strongly Concealed Stego. arXiv:2406.04218, 2024.
A.5 Example of cross-modal secret information hiding
Table 14: Examples of longer stegos generated by DAIRstega using cross-modal information. Bitstream is the binary corresponding to secret.
Secret&Bitsream Instruction Stegos PPL $\Delta$LDA
Secret:
Bitstream:
100010010101 000001001110 010001110000 110100001010 000110100000 101000000000 $000$ 000000001101 010010010100 100001000100 010100100000 $000$ 000000000110 110000000000 00…(31520 bit)
Discuss the
causes of the
Great Depression.
The Great Depression was a severe worldwide economic depression that took place from 1929 to 1939. It was the longest and most severe depression ever experienced by the industrialized world, surpassing in overall duration the Great Depression of 1873. The Depression began in the United States after the stock market crash of October 1929, and peaked between 1930 and 1933. By some measures, the Great Depression was the longest economic depression in the history of the industrialized world. The Great Depression had many causes, including overproduction, lack of consumer demand, high tariffs, banking crisis, stock market crash, deflation, wage reductions, and the 1929 collapse of Worldwide prices for agricultural commodities. The collapse of agricultural prices resulted from overproduction, due to improvements in seed stock, fertilizers, pesticides, transportation, and farm equipment, as well as increases in acreage and livestock numbers.
2.5611
0.021
Write a general overview of quantum computing.
Quantum computing is a new and revolutionary form of computing that is based on the principles of quantum mechanics. Unlike traditional computing, which uses binary bits to represent information, quantum computing uses quantum bits (qubits) which can exist in a superposition of 0 and 1 simultaneously. This means that quantum computers can process information exponentially faster than traditional computers, and they can also solve certain types of problems that are intractable for traditional computers. Quantum computing has the potential to revolutionize many fields, including machine learning, cryptography, drug discovery, and weather forecasting. For example, quantum computers can be used to train machine learning models more efficiently, which could lead to more accurate and efficient models. They can also be used to break current encryption methods, which could have serious implications for cybersecurity. In the field of drug discovery, quantum computers can be used to simulate molecular interactions, which could lead to the discovery of new and more effective drugs. Finally, in the field of weather forecasting, quantum computers can be used to simulate the complex interactions between different weather systems, which could lead to more accurate and precise forecasts.
1.9153
0.009

### v5

@@
-DAIRstega: Dynamically Allocated Interval-Based Generative Linguistic Steganography with Roulette Wheel
+Dynamically Allocated Interval-Based Generative Linguistic Steganography with Roulette Wheel
 Abstract
-Linguistic steganography (LS) tasks aim to generate steganographic text (stego) based on secret.
-Only authorized receivers can perceive and extract secrets, thereby protecting privacy.
-However, existing generative LS schemes often do not consider the conditional probability of tokens in the candidate pool, and allocate one or the same number of codings to all tokens.
-The tokens with lower probabilities are selected to embed secrets that will affect the quality of stegos.
-As a result, the stegos are easy to perceive and detect.
-This paper proposes the LS scheme based on dynamically allocated intervals, called DAIRstega.
-DAIRstega uses the idea of the roulette wheel and takes the conditional probabilities of tokens as the main basis for allocating the roulette area (i.e., the interval length).
-Thus, the token with a larger conditional probability is allocated more.
-The secret will be more likely to select the tokens with larger probabilities.
-In the allocation process, we design some functions between probability and allocated interval length.
-Based on the invisible characteristics of LS, we give three constraints that need to be met to design the function.
-To simplify the form, the expression of the allocation function is condensed.
-Furthermore, DAIRstega can receive additional instruction and controllably generate stegos.
-Rich experiments show that the proposed embedding way and DAIRstega perform superior to the existing ways and LS schemes, which shows strong perceptual, statistical, and semantic concealment and anti-steganalysis ability.
-This scheme can also generate high-quality longer stegos, improving the deficiencies in this task.
-The experiment also verified that DAIRstega can be used as a secure watermarking scheme, providing some ideas for its development.
+Existing linguistic steganography schemes often overlook the conditional probability (CP) of tokens in the candidate pool, allocating the one coding to all tokens, which results in identical selection likelihoods.
+This approach leads to the selection of low-CP tokens, degrading the quality of stegos and making them more detectable.
+This paper proposes a scheme based on the interval allocated, called DAIRstega.
+DAIRstega first uses a portion of the read secret to build the roulette area.
+Then, this scheme uses the idea of the roulette wheel and takes the CPs of tokens as the main basis for allocating the roulette area (i.e., the interval length).
+Thus, tokens with larger CPs are allocated more area.
+The secret will have an increased likelihood of selecting a token with a higher CP.
+During allocation, we designed some allocation functions and three constraints to optimize the process.
+Additionally, DAIRstega supports prompt-based controllable generation of stegos.
+Rich experiments show that the proposed embedding way and DAIRstega perform better than the existing ways and baselines, which shows strong perceptual, statistical, and semantic concealment, as well as anti-steganalysis ability.
+It can also generate high-quality longer stegos, addressing the deficiencies in this task.
+DAIRstega is confirmed to have potential as a secure watermarking, offering insights for its development.
+Our codes and data are available at: https://github.com/WangYH-BUPT/DAIRstega.
+keywords:
+Linguistic steganography , Embedding way , Allocation funtion , Interval allocated , Roulette wheel , Large language models , Concealment
+t1t1footnotetext: The paper has been accepted by Applied Soft Computing.t1t1footnotetext: Thanks for the support provided by MindSpore Community.t2t2footnotetext: This work is supported by the National Natural Science Foundation of China (Grant U21B2020) and supported by BUPT Excellent Ph.D.
+Students Foundation (Grant CX2023120, CX20241055).
 1 Introduction
-Steganography, a technology of information hiding (Shannon 1949), embeds secrets within media such as texts (Yang et al.
-2023, 2021) and images, obtaining steganographic media that is sensory indistinguishable from normal media.
-Only authorized receivers can perceive whether the media is steganographic and accurately extracts the secret.
-Benefiting from the ability of text to be transmitted losslessly over public channels, research on Linguistic Steganography (LS) has explosive growth (Lu et al.
-2023; Wang et al.
-2023a; Yang et al.
-2023).
-To increase the anti-detection ability of steganographic text (stego), the focus of work has shifted from modified LS (Huo and Xiao 2016) represented by synonym replacement (Kim, Zaiane, and Goebel 2010) to generative LS based on secret (Li et al.
-2021; Yang et al.
-2021).
+Steganography, a technology of information hiding [1], embeds secrets within media such as texts [2, 3] and images, obtaining steganographic media that is perceptually indistinguishable from normal media.
+Only authorized receivers can perceive whether the media is steganographic and accurately extract the secrets, thereby protecting privacy.
+Benefiting from the ability of text to be transmitted losslessly over public channels, research on linguistic steganography has explosive growth [4, 5, 2].
+To increase the anti-detection of steganographic text (stego), the focus of work has shifted from modified schemes [6] represented by synonym replacement [7] to generative schemes [8, 3].
 The concealment of stego determines the success of covert communication to a great extent.
@@
 In this way, stegos are difficult to perceive and detect by unauthorized receivers.
-Most existing LS schemes have made efforts in this regard.
+Most existing schemes have made efforts in this regard.
 Yang et al.
-(Yang et al.
-2019) and Ding et al.
-(Ding et al.
-2023a) encoded each token in the candidate pool with Fixed length coding (FLC) and only use secret information to determine the next token to achieve the purpose of embedding.
+[9] and Ding et al.
+[10] encoded each token in the candidate pool with Fixed length coding (FLC) and only used secret information to determine the next token to achieve the purpose of embedding.
 Yang et al.
-(Yang et al.
-2021) and Zhou et al.
-(Zhou et al.
-2021) designed the Huffman coding-based (HC) embedding, which generates stegos with good completeness and fluentness, improving perceptual concealment.
+[3] and Zhou et al.
+[11] designed the Huffman coding-based (HC) embedding, which generates stegos with good completeness and fluency, improving perceptual concealment.
 Zhang et al.
-(Zhang et al.
-2021) and Wang et al.
-(Wang et al.
-2023a) proposed the Adaptive dynamic grouping coding-based (ADG) embedding.
+[12] proposed the Adaptive dynamic grouping coding-based (ADG) embedding.
 They achieved strong statistical concealment.
+Kaptchuk et al.
+[13] and Wang et al.
+[5] proposed the Arithmetic coding-based (AC) embedding.
 Lu et al.
-(Lu et al.
-2023) used contextual learning and variable length coding-based (VLC) to generate stegos.
+[4] used contextual learning and variable length coding-based (VLC) to generate stegos.
 It improves semantic concealment.
 However, the embedding ways of almost all existing schemes allocate one or the same number of codings to the tokens in the candidate pool.
-That is, when embedding, the probability of selecting all tokens is the same.
+That is, when embedding, the likelihood of selecting all tokens is the same.
 They do not consider that the CP of a portion of the tokens is not high.
 If the secret is determined to these tokens, the quality of stego generation will be dropped.
-This will affect the quality and statistical distribution, and thus the anti-detection ability.
-The results of the existing embedding ways are given in Figure 1.
+The example is shown in the Figure 1.
+Figure 1: Examples of existing embedding ways and the proposed embedding way.
+After selecting $n$ tokens with larger CPs to build a candidate pool, the existing way assigns a coding to each token in the candidate pool.
+Since the secret binary can be arbitrary, to embed the secret, the existing embedding way selects all tokens with the same likelihood of $1/n$.
+The proposed way will build a roulette area and allocate it according to the CP value to ensure that the larger the CP of the token, the larger the allocation area.
+Therefore, the proposed way has a higher likelihood of selecting a better token.
+In the end, the quality of the token determined by the proposed way is generally better than that of the existing ways.
+Tokens with low CPs will reduce the quality of stego and cause significant differences in its statistical characteristics from cover.
+Since steganalysis detects stegos by capturing these differences [14, 15], poor-quality stego is easier to identify and has weaker anti-steganalysis capabilities.
+The performance of stegos obtained by different embedding ways is shown in Figure 2.
 Therefore, it is necessary to explore a difficult-to-perceive and detect scheme with concealed embedding.
@@
 $\uparrow$” and $\downarrow$” represent the higher / lower the value, the better the result.
-The metrics can be found in “Section 3.1”.
-The values can be seen in Table 1.
-This paper proposes the dynamically allocated interval-based LS scheme, called DAIRstega.
-Different from the existing LS, DAIRstega introduces the idea of the roulette wheel to design the secret embedding way.
+The metrics are found in “Section 4.1”.
+The values are seen in Table 2.
+This paper proposes the dynamically allocated interval-based steganography, called DAIRstega.
+Different from the existing schemes, DAIRstega introduces the idea of the roulette wheel to design the embedding way.
 We believe that the quality of the tokens in the candidate pool is different.
-Thus, the tokens need to be given different shoot probabilities by secret, making it more likely that the secret will select tokens with higher CPs.
-Then, we explore the characteristics of the CP of tokens in the candidate pool, and design a variety of relationships between CPs and the number of allocated codings.
-By analyzing the adaptability of these allocated functions to the invisible information-hiding tasks of LS, the allocation expression is condensed.
-In addition, DAIRstega can receive instruction and more directly control the generation of the stego contents.
-Experiments show that DAIRstega has strong concealment and controllability.
-The main contributions are summarized as follows.
+Thus, the tokens need to be given different shoot likelihoods by secret, increasing the likelihood of selecting tokens with higher CP.
+Then, we explore the characteristics of the CP of tokens, and design a variety of relationships between CPs and the number of allocated codings.
+By analyzing the adaptability of these allocated functions to the invisible information-hiding tasks, the allocation expression is condensed.
+In addition, DAIRstega can receive instructions and directly control the stego generation.
+DAIRstega is built upon the MindSpore framework and the contributions are summarized as follows.
+  • 1.
     To augment the concealment of stegos, DAIRstega introduces the idea of the roulette wheel, and dynamically allocates the number of codings.
-The multiple functions between CPs and allocated numbers are also given.
+The functions between CPs and allocated numbers are also given.
 According to Occam’s razor, the functions are condensed.
-  • To ensure the adaptability of the functions designed, we gave three constraints that need to be met.
+  • 2.
+    To ensure the adaptability of the functions designed, we gave three constraints.
+  • 3.
     To make the evaluation results consistent with the LLMs era, the high-quality corpus was built and used this corpus as a cover to train the model of all schemes.
-  • To verify the effectiveness of the proposed embedding way and DAIRstega, we performed experiments regarding various concealment and anti-steganalysis.
-  • Experiment verified that DAIRstega can also be regarded as a secure watermarking scheme with high fidelity, consistency, and abuse-resistant.
+  • 4.
+    To verify the effectiveness of the proposed embedding way and DAIRstega, we performed experiments regarding perceptual, statistical, and semantic concealment, as well as anti-steganalysis performance.
+  • 5.
+    Experiment verified that DAIRstega can also be regarded as a secure watermarking scheme with high quality, consistency, and unforgeability.
+2 Related work
+The concealment of stego determines the success of covert communication to a great extent. The concealment is primarily manifested in three aspects: perceptual, statistical, and semantic.
+Perceptual concealment” focuses on ensuring that the scheme generates stego with complete and fluent.
+This is the most basic requirement of steganography.
+In pursuit of this objective, Fang et al.
+[16] proposed an LSTM-based steganography scheme.
+This scheme segments the vocabulary into several sets based on bit blocks.
+It then selects tokens with the highest probability that corresponds to the secret information from the candidate pool.
+Ding et al.
+[10] combined the conditional generation strategy with the replacement technique, using text sequences as auxiliary data in the stego-generation process to enhance the embedding capabilities.
+Yang et al.
+[9] trained a language model to generate stegos.
+Using HC-based (Huffman coding) and FLC-based (Fixed length coding) to encode the candidate pool, these stegos have excellent completeness and fluentness.
+Statistical concealment” requires the stegos distribution to be close to the covers.
+To achieve this goal, Yang et al.
+[3] designed an encoder-decoder structure, and VLC-based (Variable length coding) is used to encode the CP.
+To further reduce the distribution difference between cover and stego, Zhang et al.
+[12] used adaptive dynamic grouping coding to embed the secret recursively.
+Zhou et al.
+[11] used a generative adversarial network (GAN) to design an adaptive probability distribution steganography.
+These ensure statistical concealment.
+Semantic concealment” aims to generate a stego that is coherent and semantically controllable.
+To this end, Li et al.
+[8] put forward a steganography scheme based on the knowledge graph.
+Yang et al.
+[2] utilized semantic information encoding to embed secret information, realizing the effect of maintaining semantics and increasing the embedding capacity during the translation process.
+Wang et al.
+[5] leveraged the relevance of social network context to enhance contextual semantic relevance while maintaining embedding rates.
+Lu et al.
+[4] used contextual learning and GPT2 guidance to generate stegos.
 3 Methodology
@@
 Figure 3: Example of texts being transmitted and detected. Responses that match the discourse are difficult to perceive, thus reducing the risk of steganalysis, such as Alice 1.
-The covert communication hopes to avoid Eve’s perception and detection and needs to satisfy the following:
+The covert communication hopes to avoid Eve’s perception and detection and needs to satisfy:
 $\left\{\begin{array}[]{l}d({P_{C}},{P_{S}})\leq\varepsilon\\ C\Rightarrow\mathcal{C}_{style}\sim{P_{\mathcal{C}1}},{S_{k}}\Rightarrow\mathcal{S}_{style}\sim{P_{\mathcal{S}1}}\\ C\Rightarrow\mathcal{C}_{genre}\sim{P_{\mathcal{C}2}},{\rm{}}{S_{k}}\Rightarrow\mathcal{S}_{genre}\sim{P_{\mathcal{S}2}}\\ (d({P_{\mathcal{C}1}},{P_{\mathcal{S}1}})\leq{\varepsilon_{1}})\wedge(d({P_{\mathcal{C}2}},{P_{\mathcal{S}2}})\leq{\varepsilon_{2}})\end{array}\right.,$ (1)
@@
 Complying with the Formula 1 can improve the concealment and ensure the conduct of covert communication.
-The results of this aspect are shown in Table 4 and Table 5.
+The results of this aspect are shown in Table 5.
 3.2 DAIRstega Overall
 This scheme can generate stegos that are strongly concealed and controllable.
-It can satisfy the extension of covert communication, i.e.
-Formula 1.
-To make the generated stegos more consistent with the CP distribution of covers, DAIRstega uses the idea of the roulette wheel to dynamically allocate tokens in the candidate pool to coding intervals of different lengths.
+It can satisfy the extension of covert communication, i.e., Formula 1.
+To make the generated stegos more consistent with the CP distribution of covers, DAIRstega uses the idea of the roulette wheel to allocate tokens dynamically in the candidate pool to coding intervals of different lengths.
 That is, the larger the CP is, the larger the roulette area will be allocated, and the greater the shoot that this token will be selected by secret.
-To make the generated stegos controllable, we introduce the Prompt, taking instructions to control generate.
-Instruction can be public topics, posted questions, or other requirements in social networks.
+Instructions can be public topics, posted questions, or other requirements.
 Figure 4 shows the DAIRstega’s framework.
 Figure 4: The DAIRstega’s framework.
-The “stego generation process” consists of two parts: the “CP generation” and “Embedding” modules.
-Different from existing LS works, DAIRstega uses the idea of the non-uniform roulette wheel to dynamically allocate different roulette areas (numbers of codings) to tokens in the candidate pool.
+The “stego generation process” consists of two parts: the “CP (conditional probability) generation” and “Embedding” modules.
+Different from existing works, DAIRstega uses the idea of the non-uniform roulette wheel to dynamically allocate different roulette areas (numbers of codings) to tokens in the candidate pool.
 In this way, the secret is more likely to fall on the tokens with larger CPs, ensuring the quality of stegos.
@@
 3.3 Details
-Stego generation process.
-It is composed of the “CP generation” ($PG$) and “Embedding” ($Em$) modules. The process formulas are as follows:
-$\left\{\begin{array}[]{l}PG:LM(I)\Rightarrow\mathbf{P}\\ Em:\text{locate}\;\{{m_{t}}\;\text{in}\;DAI(BC(\mathbf{P}))\}\Rightarrow i,\\ \quad\quad\;\ \text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}\\ i=Em(PG(I))\end{array}\right.,$ (2)
+Stego generation process
+It is composed of the “CP (conditional probability) generation” ($CPG$) and “Embedding” ($Em$) modules. The process formulas are as follows:
+$\left\{\begin{array}[]{l}CPG:LM(I)\Rightarrow\mathbf{P}\\ Em:\text{locate}\;\{{m_{t}}\;\text{in}\;DAI(BC(\mathbf{P}))\}\Rightarrow i,\\ \quad\quad\;\ \text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}\\ i=Em(CPG(I))\end{array}\right.,$ (2)
 where, $I$ and $\mathbf{P}$ are the index sequence of a given text and the CP distribution of the next token, $\mathbf{P}=\left[{p_{1}},\cdots,{p_{v}}\right]$, ${p_{v}}$ is the CP of the $v$-th token in the vocab.
@@
 “$\text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}$” is the same prefix that matches the beginning and end of $i$’s range.
-CP generation module.
+CP generation module
 The input is the instruction or the instruction with part of the subsequent token generated, and the output is the CP distribution $\mathbf{P}$ of the next token.
-Embedding module.
-To ensure that the embedding does not excessively affect the quality of the generated stegos.
+Embedding module
+To ensure that the embedding does not excessively affect the stego quality.
 In our work, low-CP tokens should not be added to the candidate pool.
@@
 […]
-Further exploration and discovery, if the linear function is used, the embedding rate can be improved.
-Tokens with very high $p$ will be allocated to most encodings, and a large interval length will result in a short prefix of this interval regardless of how many bits are read at a time.
-Longer stego needs to be generated to embed the given bits, resulting in a lower embedding rate (unit: bit per word).
-So we try to design the nonlinear functions to improve the embedding rate.
+Further exploration and discovery, if the linear function is used, the payload (unit: bit per word) can be improved.
+Tokens with very high $p$ will be allocated to most codings, and a large interval length will result in a short prefix of this interval regardless of how many bits are read at a time.
+Longer stego needs to be generated to embed the given bits, resulting in a lower payload.
+So, we try to design the nonlinear functions to improve the payload.
 Using logarithmic and exponential operations, a variety of nonlinear allocations can be designed.
-However, not all nonlinear allocations are suitable for LS information hiding.
-We define three constraints that the functions need to be met:
-Constraint 1:
-${\frac{{\partial\ {DAI}{(x)}}}{{\partial\ x}}}\geq\frac{c}{1-x},c\in\mathbb{R}^{+},\forall x\in\left[\epsilon^{+},1\right),$ (5)
-where, $x$ is a variable that meets all CPs.
-$\epsilon^{+}$ is a smaller value greater than 0, such as 0.1, 0.2, and so on.
+However, not all nonlinear allocations are suitable for information hiding.
+We define three constraints that the functions need to meet.
+$x$ is a variable that meets all CPs.
+Since the CP value of the token in the candidate pool is relatively large, $\epsilon^{+}$ is a smaller value greater than 0, such as 0.1, 0.2, and so on.
 Although the CP distribution is discrete, the independent variable of the distribution function is continuous, so it is expressed here using derivatives.
-This constraint shows that the number of allocations needs to increase with the increase of CP, and ensures that the change rate of the allocated number is not too small.
-If ${\textstyle{{\partial\;DAI(x)}\over{\partial\;x}}}\leq\varepsilon$, which means the number of allocation tokens is similar.
-It will degenerate into a similar allocation of the same number of codings as the existing work, thus increasing the selection of a poor token.
-Constraint 2:
-${\frac{{\partial^{2}{DAI}{(x)}}}{{\partial x^{2}}}}\leq 0,\forall x\in\left[\epsilon^{+},1\right).$ (6)
-When the condition is equal, the corresponding function will be transformed into a linear function.
-This constraint indicates that the entire allocation function should be an increasing function with decreasing growth.
-If the function has an increasing growth trend, then tokens with larger CPs will be allocated to larger intervals, and the same prefixes at the beginning and end of the interval will be shortened.
-Thus, the embedding rate will be decreased.
-Constraint 3:
-$DAI(x)\geq x\;\wedge\;{\frac{{\partial\ {DAI}{(x)}}}{{\partial\ x}}}-{\frac{{\partial\ x}}{{\partial\ x}}}\leq 0,\forall x\in\left[\epsilon^{+},1\right).$ (7)
-This constraint shows that the number of codings allocated to tokens with smaller CP should be appropriately increased to increase their probability of being selected.
-This can also suppress the problem of too large intervals of larger CP to a certain extent.
-Here, we give several nonlinear allocation functions:
+Constraint 1: (Guaranteed text diversity)
+$DAI(x)\geq x,\forall x\in\left[\epsilon^{+},1\right).$ (5)
+If the steganographic embedding can only select a token with extremely high CP, the generated diversity will be lost, and the existence of the stegos will be easily exposed.
+This constraint indicates that the number of codings allocated to tokens with lower CPs should be appropriately increased to enhance their selection likelihood.
+Constraint 2: (Guaranteed text quality)
+${\frac{{\partial\ {DAI}{(x)}}}{{\partial\ x}}}\geq\epsilon^{+},\forall x\in\left[\epsilon^{+},1\right),$ (6)
+The existing embedding ways assign the same number of codings to all tokens, which results in the same likelihood of randomly selecting each token, thus affecting the generation quality.
+Therefore, this constraint requires that tokens with larger CP need to be assigned a longer interval, that is, the number of allocations needs to increase with the increase of CP, and the change rate of the number of allocations cannot be too small.
+If ${\textstyle{{\partial\;DAI(x)}\over{\partial\;x}}}\leq\epsilon^{+}$, which means the number of allocation tokens is similar.
+It will degenerate into a similar allocation of the same number of codings as the existing work, thereby increasing the likelihood of selecting poor tokens.
+Constraint 3: (Guaranteed to facilitate embedding)
+${\frac{{\partial\ {DAI}{(x)}}}{{\partial\ x}}}\leq{\frac{{\partial\ x}}{{\partial\ x}}}\cap{\frac{{\partial^{2}{DAI}{(x)}}}{{\partial x^{2}}}}\leq 0,\forall x\in\left[\epsilon^{+},1\right).$ (7)
+Steganography requires an algorithm that is conducive to embedding information.
+If you just want high-quality tokens to be selected and blindly increase the interval length of tokens with high CPs, the same prefix at the beginning and end of the interval will be shortened or even non-existent.
+This is not conducive to embedding secrets and deviates from the basic function of steganography.
+Therefore, in order to ensure the payload, it is necessary to avoid high CP tokens occupying too much interval length, and the growth trend of ${DAI}{(x)}$ should be gradually slow, that is, mathematically, it can be achieved by limiting the second-order derivative to be less than 0.
+When the conditions are equal, the corresponding function will be transformed into a linear function.
+Here, we give several nonlinear functions:
 […]
@@
 The smaller $\beta$ is, the smaller the interval of tokens with larger CP will be, while still satisfying the premise that the interval of tokens with larger CP is larger.
-It will increase the length of the prefix, while reducing the length of stegos to be generated, improving the embedding rate.
+It will increase the length of the prefix, while reducing the length of stegos to be generated, improving the payload.
 Please note that these functions are not fixed, here are just a few examples that meet the constraints.
-The constraints we proposed are not the most perfect, and the final condensed expression is also for simplicity and easy to understand, and is not the best.
+These constraints are not exhaustive, and the final condensed expression is also for simplicity and easy to understand.
 This can just provide thinking for subsequent research that is more conducive to hiding information.
-Loop the stego generation process, the stego will be obtained. Algorithm 1 illustrates the embedding process.
+Loop the stego generation, and stego will be obtained. Algorithm 1 illustrates the embedding process.
 Algorithm 1 Secret embedding of DAIRstega.
@@
 14:  Map $[{i_{n+1}},\cdots,{i_{{\rm{<EOS>}}}}]\Rightarrow[{t_{n+1}},\cdots,{\rm{<EOS>}}]$.
-Secret extraction process.
+Secret extraction process
 Bitstream extraction and embedding are a pair of inverse operations.
-Same as Discop (Ding et al.
-2023b), we also need to use the same scheme, initial context (instruction), and stegos.
+Same as Discop [17], we also need to use the same scheme, initial context (instruction), and stegos.
 The instructions remain requirements visible to everyone in social networks.
 The formula of the extraction process $Ex$ is as follows:
-$\left\{\begin{array}[]{l}Ex:\text{locate}\;\ \{{i_{n+1}}\;\ {\rm{in}}\;\ DAI(BC(P))\},\\ \quad\quad\;\text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}\\ B_{j}=Ex(PG(I)),B_{j}\in B\end{array}\right..$ (10)
-The interval is located by the stego’s next token at this moment, and the same prefix is the current extracted secret.
+$\left\{\begin{array}[]{l}Ex:\text{locate}\;\ \{{i_{n+1}}\;\ {\rm{in}}\;\ DAI(BC(P))\},\\ \quad\quad\;\text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}\\ B_{j}=Ex(CPG(I)),B_{j}\in B\end{array}\right..$ (10)
+The interval is located by the stego’s next token at this moment, and the same prefix is the currently extracted secret.
 4 Experiments
@@
 4.1 Settings
-Dataset.
-We obtained a large amount of data from Wikipedia, Twitter, GPT4 (OpenAI 2023), and the publicly available data for training LLMs111 https://github.com/tloen/alpaca-lora/tree/main.
-The final dataset contains 108,771 texts and nearly 100 different discourse characteristics to train the language models of all schemes.
-Model configuration.
+Dataset
+We obtained a large amount of data from Wikipedia, Twitter, GPT4 [18], and the publicly available data for fine-tuning LLMs111 https://github.com/tloen/alpaca-lora/tree/main.
+The final dataset contains 57,414 texts and nearly 100 different discourse characteristics to fine-tune the language models of all schemes.
+MindSpore framework
+All experiments are based on the MindSpore framework.
+MindSpore is an open-source AI framework introduced by Huawei, designed to support flexible deployment across cloud, edge, and device scenarios.
+In this research, MindSpore played a significant role in unleashing the computational power of the hardware.
+Model configuration
+LLM is divided into closed-source and open-source LLMs, closed source such as Gemini, GPT4, etc., open-source such as LLaMA2-7B, etc.
+For detailed information on closed-source LLMs, please refer to the review of comprehensive research in related fields [19, 20].
+Since DAIRstega needs to modify the token sampling inside LLMs, this paper uses open-source LLMs for experiments.
 DAIRstega can access any open-source LLMs.
-In the experiment, we used LLaMA2-7B and 13B models (Meta 2023) as an example.
-The rank of LoRA (Hu et al.
-2021) is 8, and the dropout is 0.05.
-The learning algorithm is AdamW (Loshchilov and Hutter 2018), and the learning rate is 1e-5.
-The batch size is 16.
+In the comparison experiments, the LLaMA2-7B model is used [21] as the base LLM.
+The LLaMA2-7B and LLaMA2-13B models are used in the ablation experiments of the different LLMs, and all models are operated with 16-bit precision.
 For simplicity, we use $\beta=1,0.5$ and $\alpha=8,32,48$ to perform DAIRstega.
-Table 1: The overall results of different embedding ways.
+Baselines
+Since the steganography scheme consists of a language model and an embedding way, the language models of existing steganography are different.
+Therefore, to ensure fairness in the comparison, we not only compared the schemes, but also compared the performance of different embedding ways.
+That is to say, comparison is divided into two parts: one is the comparison between the proposed embedding way and the existing embedding ways, and the other is the comparison between DAIRstega and the prevalent works in recent years.
+Table 1: Overview of each part of the experiments. It includes: experimental part (Focus), Section, Metrics, Base LLM, Table, and supplementary materials for table (Supplement).
+Focus Section Metrics Base LLM Table Supplement
+Comparison study Embedding ways 4.2 Perceptual, Statistical, Semantic, Anti-steganalysis LLaMA2-7B 2 10, 11
+Steganography baselines 4.3.1 Perceptual 3
+4.3.2 Statistical 4
+4.3.3 Semantic 5 12
+4.3.4 Anti-steganalysis 6 13, 7.3
+Related-task baseline 4.4 Quality, Consistency, Unforgeability 7
+DAIRstega exploration Longer stegos 4.5 Perceptual, Semantic LLaMA2-7B 8 15
+Ablation of LLMs 4.6 Perceptual, Statistical, Semantic, Anti-steganalysis LLaMA2-7B, 13B 9
+Table 2: The overall results of different embedding ways (based on MindSpore framework).
 The embedding rate of each way is near 1Bpw (bit per word).
 Bold and “ *” are the best and the suboptimal result.
-“LLM” in this table and Ours are all selected the LLaMA2-7B.
-For the LS works involved in these embedding ways, see “Section 3.1”.
+The “LLM”s in this table are all LLaMA2-7B.
+For the steganography works involved in these embedding ways, see “Section 4.1”.
 $\uparrow$” is that the higher the value, the better the result.
 $\downarrow$” means the lower the value, the better the result.
-The complete data can be seen in Table 8 and Table 9 in Supplementary A.2.
+The complete data can be seen in Table 10 and Table 11 in Appendix 7.1.
 […]
@@
 LLM+FLC 12.883 5.190 97.04 46.56 0.241 8.280 2.96 79.48 54.88
-Ours 8.030 0.342 97.43 38.60 0.227 6.341 2.57 78.53* 77.42
-Embedding $\hookrightarrow$ Anti-steganalysis
-Score $\uparrow$ LS_CNN TS_CSW EILG UP4LS
+LLM+Ours 8.030 0.342 97.43 38.60 0.227 6.341 2.57 78.53* 77.42
+Embedding ways $\hookrightarrow$ Anti-steganalysis
+Score $\uparrow$ LS_CNN [30] TS_CSW [31] EILG [32] UP4LS [33]
 Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$
@@
 LLM+FLC 58.44 78.53 80.05 82.72 83.32 76.67 76.15 89.03 89.20
-Ours 65.41 58.95 58.74 55.28 61.69 65.01 65.83 72.92 73.30
-Baselines.
-Comparison is divided into two parts, one is the comparison between the proposed embedding way and the existing embedding ways, and the other is the comparison between DAIRstega and the prevalent works in recent years.
-$\cdot$ In terms of the embedding ways, the existing LS works can be summarized as follows: (1) HC (or VLC), (2) ADG, (3) AC, and (4) FLC.
-(1) HC is used in VAEstega (Yang et al.
-2021), LSCS (Lu et al.
-2023), Hi-Stega (Wang et al.
-2023a), RNNstega (Yang et al.
-2019), GAN-APD (Zhou et al.
-2021), JLS (Ding et al.
-2023a), Discop (Ding et al.
-2023b).
-(2) ADG is used in Hi-Stega (Wang et al.
-2023a), ADG (Zhang et al.
-2021).
-(3) AC is used in VAEstega (Yang et al.
-2021), Hi-Stega (Wang et al.
-2023a).
-(4) FLC is used in RNNstega (Yang et al.
-2019), JLS (Ding et al.
-2023a).
-$\cdot$ In terms of the prevalent works, we selected the LS and watermarking works that showed excellent performance as the baselines.
-LS: (1) Fang (ACL) (Fang, Jaggi, and Argyraki 2017).
-(2) RNNstega (IEEE TIFS) (Yang et al.
-2019).
-(3) ADG (ACL-IJCNLP) (Zhang et al.
-2021).
-(4) LSCS (IJCNN) (Lu et al.
-2023).
-Watermarking: (5) PLMmark (AAAI) (Liu et al.
-2023).
-It uses pre-trained language models to provide the first secure black-box watermarking.
+LLM+Ours 65.41 58.95 58.74 55.28 61.69 65.01 65.83 72.92 73.30
+  • 1. HC (Huffman coding) is used in VAE-Stega [3], LSCS [4], RNNstega [9], GAN-APD [11], JLS [10], and Discop [17].
+  • 2. ADG (Adaptive dynamic grouping) is used in Hi-Stega [5] and ADG [12].
+  • 3. AC (Arithmetic coding) is used in VAE-Stega [3], Hi-Stega [5], and Meteor [13].
+  • 4. FLC (Fixed length coding) is used in RNNstega [9] and JLS [10].
+$\cdot$ In terms of the embedding ways, the existing works can be summarized as follows: (1) Huffman coding (HC) (or Variable length coding (VLC)), (2) Adaptive dynamic grouping (ADG), (3) Arithmetic coding (AC), and (4) Fixed length coding (FLC).
+(1) HC is used in VAE-Stega [3], LSCS [4], RNNstega [9], GAN-APD [11], JLS [10], Discop [17].
+(2) ADG is used in Hi-Stega [5] and ADG [12].
+(3) AC is used in VAE-Stega [3], Hi-Stega [5], and Meteor [13].
+(4) FLC is used in RNNstega [9], JLS [10].
+$\cdot$ In terms of the prevalent works, we selected the steganography and watermarking works that showed excellent performance as the baselines.
+Steganography-task baselines: (1) Fang [16].
+(2) RNNstega [9].
+(3) ADG [12].
+(4) LSCS [4].
+Related-task baseline: (5) PLMmark [22].
+It uses pre-trained language models to provide the first secure black-box watermarking scheme.
 They can compare the generation effects of DAIRstega from different aspects, and these performances have been widely recognized.
-Evaluation metrics.
+Evaluation metrics
 $\downarrow$ means the lower the value, the better, $\uparrow$ means the higher the value, the better.
-$\cdot$ In terms of text quality (perceptual concealment), we adopt (1) PPL $\downarrow$ (stego perplexity).
-(Wu et al.
-2023; Ding et al.
-2023a).
-(2) $\Delta$Pcs $\downarrow$ (The difference between PPL of cover and stego (Yang et al.
-2021)).
-$\cdot$ In terms of statistical analysis (statistical concealment), we adopt (3) CS $\uparrow$.
+$\cdot$ In terms of perceptual concealment (text quality), we adopt (1) PPL $\downarrow$ (stego perplexity). [23, 10]. (2) $\Delta$Pcs $\downarrow$ (The difference between PPL of cover and stego [3]).
+$\cdot$ In terms of statistical concealment (statistical analysis), we adopt (3) CS $\uparrow$.
 (Cosine similarity).
@@
 (6) MD $\downarrow$ (Manhattan distance).
-(7) $\Delta$DP $\downarrow$ (Dot product difference (Cai et al.
-2021)).
-$\cdot$ In terms of discourse matching (Semantic concealment), we adopt (8) $\Delta$LDA $\downarrow$.
-(9) Mauve $\uparrow$ (Pillutla et al.
-2021).
+(7) $\Delta$DP $\downarrow$ (Dot product difference [24]).
+$\cdot$ In terms of semantic concealment (discourse matching), we adopt (8) $\Delta$LDA $\downarrow$.
+(9) Mauve $\uparrow$ [25].
 Mauve is a metric that considers fluency and consistency.
-(10) BLEU $\uparrow$ (Bilingual evaluation understudy (Papineni et al.
-2002)).
-(11) Score (BERTScore) $\uparrow$ (Zhang et al.
-2020).
+(10) BLEU $\uparrow$ (Bilingual evaluation understudy [26]).
+(11) Score (BERTScore) $\uparrow$ [27].
 Score considers the similarity at the segment level.
-$\cdot$ In terms of anti-steganalysis, we use the high-performance linguistic steganalysis LS_CNN (Wen et al.
-2019), TS_CSW (Yang, Huang, and Zhang 2020), EILG (Xu, Zhang, and Liu 2023), and UP4LS (Wang et al.
-2023b): (12) Acc $\downarrow$ (Accuracy) and (13) F1 $\downarrow$ (F1-score).
-3.2 Comparison of the embedding ways
-To illustrate the effect of embedding way, we performed this experiment. We add the existing works with the same LLMs as ours. The results are shown in Table 1.
-The results in Table 1 show that the proposed embedding way can improve the stego quality, better simulate the CP distribution of covers, and get the optimal anti-steganalysis.
+$\cdot$
+In terms of anti-steganalysis, steganalysis is essentially a detection method [28].
+It can detect whether there is secret information in the text [29].
+we use the high-performance linguistic steganalysis LS_CNN [30], TS_CSW [31], EILG [32], and UP4LS [33]: (12) Acc $\downarrow$ (Accuracy) and (13) F1 $\downarrow$ (F1-score).
+Experimental overview
+We provide an overview of each part of the experiments, as shown in Table 1.
+4.2 Comparison with the different embedding ways
+To illustrate the effect of the embedding way, we performed this experiment. We add the existing works with the same LLMs as ours. The results are shown in Table 2.
+Table 2 shows that the proposed embedding way can improve the stego quality (perceptual concealment), better simulate the CP distribution of covers (statistical concealment), and get the optimal anti-steganalysis.
 Overall performance performs better than the existing ways.
-3.3 Comparison with LS baselines
-Text quality (Perceptual concealment).
-Table 2 shows the comparison regarding the stego quality.
-Table 2: Text quality comparison of stegos generated by DAIRstega and LS baselines.
+4.3 Comparison with steganography-task baselines
+4.3.1 Perceptual concealment (text quality)
+Table 3 shows the comparison regarding the perceptual concealment.
+Table 3: Perceptual concealment comparison of stegos generated by DAIRstega and steganography baselines.
 Bold is the best result.
@@
 Schemes Param / Bpw PPL $\downarrow$ $\Delta$Pcs $\downarrow$
-Fang bin = 1 / 1.00 235.271 227.578
+Fang [16] bin = 1 / 1.00 235.271 227.578
 bin = 3 / 3.00 269.303 261.611
-RNNstega bit = 1 / 1.00 79.987 72.294
+RNNstega [9] bit = 1 / 1.00 79.987 72.294
 bit = 3 / 2.63 99.302 91.610
-ADG $\tau=0.5$ / 4.38 197.498 189.806
-LSCS - / 1.12 9.248 1.556
+ADG [12] $\tau=0.5$ / 4.38 197.498 189.806
+LSCS [4] - / 1.12 9.248 1.556
 DAIRstega $\beta$ = 1 $\alpha$ = 8 / 1.10 7.686 0.006
@@
 This shows that DAIRstega can be more consistent with the linguistic characteristics of covers, enhancing the perceptual concealment of the stegos.
-Statistical analysis (Statistical concealment).
+4.3.2 Statistical concealment (statistical analysis)
 Table 4 shows the comparison between DAIRstega and baselines regarding the statistical analysis between covers and stegos.
-Table 3: Statistical analysis comparison of stegos generated by DAIRstega and LS baselines. The meanings of “Bold”, “ *”, “bin”, “bit”, and “$\tau$” are the same in Table 2.
+Table 4: Statistical concealment comparison of stegos generated by DAIRstega and steganography baselines. The meanings of “Bold”, “ *”, “bin”, “bit”, and “$\tau$” are the same in Table 3.
 Schemes Param CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$ MD $\downarrow$ $\Delta$DP $\downarrow$
-Fang bin = 1 9.31 63.43 1.347 17.911 90.69
+Fang [16] bin = 1 9.31 63.43 1.347 17.911 90.69
 bin = 3 9.30 63.34 1.347 17.876 90.70
-RNNstega bit = 1 95.97 41.51 0.284 7.530 4.03
+RNNstega [9] bit = 1 95.97 41.51 0.284 7.530 4.03
 bit = 3 96.32 39.38 0.271 7.371 3.68
-ADG $\tau=0.5$ 96.82 43.02 0.252 8.232 3.18
-LSCS - 94.20 40.28 0.340 7.182 5.80
+ADG [12] $\tau=0.5$ 96.82 43.02 0.252 8.232 3.18
+LSCS [4] - 94.20 40.28 0.340 7.182 5.80
 DAIRstega $\beta$=1 $\alpha$=8 97.20 38.86 0.237 6.458 2.80
@@
 According to the comparison results in Table 4, it can be found that DAIRstega can better simulate the statistical distribution of cover, enhancing the statistical concealment.
-Discourse matching (Semantic concealment).
-Table 4 shows the comparison regarding discourse matching.
-Table 4: Discourse-matching comparison of stegos generated by DAIRstega and LS baselines. Unit is %. The complete data can be seen as Table 11 in Supplementary A.3.
+4.3.3 Semantic concealment (discourse matching)
+Table 5 shows the comparison regarding semantic concealment.
+Table 5: Semantic concealment comparison of stegos generated by DAIRstega and steganography baselines. The unit is %. The complete data can be seen as Table 12 in Appendix 7.2.
 Schemes Param $\Delta$LDA $\downarrow$ Mauve $\uparrow$ BLEU $\uparrow$ Score $\uparrow$
-Fang bin = 1 11.964 2.15 0.90 42.58
+Fang [16] bin = 1 11.964 2.15 0.90 42.58
 bin = 3 11.964 1.96 0.90 41.65
-RNNstega bit = 1 0.062 12.73 1.87 52.57
+RNNstega [9] bit = 1 0.062 12.73 1.87 52.57
 bit = 3 18.919 11.97 1.83 52.36
-ADG $\tau=0.5$ 15.901 18.54 2.72 47.55
-LSCS - 0.022 2.40 31.95 59.22
+ADG [12] $\tau=0.5$ 15.901 18.54 2.72 47.55
+LSCS [4] - 0.022 2.40 31.95 59.22
 DAIRstega $\beta$=1 $\alpha$=8 0.017* 79.91* 76.37 65.76
@@
 $\alpha$=48 0.006 72.40 79.64 60.29
-Anti-steganalysis
+4.3.4 Anti-steganalysis
 Table 6 shows the anti-steganalysis comparison of DAIRstega and baselines.
-Table 5: Anti-steganalysis (Acc $\downarrow$) comparison of stegos generated by DAIRstega and LS baselines.
-Unit is %.
-The complete data can be seen as Table 12 in Supplementary A.3.
-We give the detection of the other 7 steganalysis (including non-BERT, BERT, and LLMs-based).
-It can be seen as Table A.4 in Supplementary A.4.
-Schemes Param LS_CNN TS_CSW EILG UP4LS
-Fang bin = 1 99.70 88.34 99.50 99.75
+Table 6: Anti-steganalysis (Acc $\downarrow$) comparison of stegos generated by DAIRstega and steganography baselines.
+The unit is %.
+The complete data can be seen as Table 13 in Appendix 7.2.
+We also give the detection of the other 7 steganalysis works [14, 34, 35, 36, 37, 38, 15] (including non-BERT, BERT, and LLM-based).
+It can be seen as Table 7.3 in Appendix 7.3.
+Schemes Param LS_CNN [30] TS_CSW [31] EILG [32] UP4LS [33]
+Fang [16] bin = 1 99.70 88.34 99.50 99.75
 bin = 3 99.75 89.03 99.52 99.88
-RNNstega bit = 1 85.81 81.39 85.86 97.42
+RNNstega [9] bit = 1 85.81 81.39 85.86 97.42
 bit = 3 82.93 80.35 82.71 96.78
-ADG $\tau=0.5$ 85.86 82.23 84.86 98.34
-LSCS - 78.86 81.19 84.15 96.22
+ADG [12] $\tau=0.5$ 85.86 82.23 84.86 98.34
+LSCS [4] - 78.86 81.19 84.15 96.22
 DAIRstega $\beta$=1 $\alpha$=8 59.50 56.82 67.25 75.40
@@
 $\alpha$=48 70.17 69.48 70.64 79.46
-3.4 Comparison with watermarking baseline
-The rise of LLMs has spawned many excellent watermarking papers (Liu et al.
-2023; Kirchenbauer et al.
-2023).
-They provide solutions for copyright protection in terms of high fidelity (Kirchenbauer et al.
-2023), consistency (Zhang et al.
-2024), and abuse resistance (Liu et al.
-2023).
-Excitingly, DAIRstega can also be regarded as an invisible watermarking with strong fidelity, consistency, and abuse resistance.
-The focus of this paper is not watermarking, we only compare it with the PLMmark (AAAI 2023), as shown in Table 6.
-Table 6: Fidelity, consistency, and abuse resistance comparison with watermarking baseline. The complete data can be seen in Table 11 and Table 12 in Supplementary A.3.
-[…]
-PPL $\downarrow$ $\Delta$Pcs $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$
-PLMmark 18.748 11.056 0.009 7.22
-DAIRstega 7.686 0.006 0.006 80.83
-Scheme $\hookrightarrow$ Abuse resistance
-Score $\uparrow$ TS_CSW $\downarrow$ EILG $\downarrow$ UP4LS $\downarrow$
-PLMmark 43.50 98.81 99.92 99.85
-DAIRstega 65.76 54.05 62.28 69.04
-According to the comparison results in Table 6, the watermark texts generated by DAIR have excellent fidelity and consistency, and it is difficult for unauthorized receivers to detect the existence of the watermark text, avoiding the problem of watermark abuse.
+4.4 Comparison with related-task baseline
+The rise of LLMs has spawned many excellent watermarking papers [22, 39].
+They provide solutions for copyright protection in terms of high text quality [39], consistency [40], and unforgeability [22].
+Excitingly, DAIRstega can also be regarded as an invisible watermarking with high quality, consistency, and unforgeability.
+The focus of this paper is not watermarking, we only compare it with the PLMmark, as shown in Table 7.
+Table 7: Quality, consistency, and unforgeability comparison with watermarking baseline. The complete data can be seen in Table 12 and Table 13 in Appendix 7.2.
+Scheme Quality Consistency Unforgeability
+PPL $\downarrow$ $\Delta$Pcs $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$ Score $\uparrow$ TS_CSW $\downarrow$ EILG $\downarrow$ UP4LS $\downarrow$
+PLMmark [22] 18.748 11.056 0.009 7.22 43.50 98.81 99.92 99.85
+DAIRstega $\beta$=1, $\alpha$=8 7.686 0.006 0.017* 79.91* 65.76 56.82 67.25 75.40
+$\beta$=1, $\alpha$=32 8.184* 0.492* 0.022 76.97 65.22 54.05 62.28 71.97
+$\beta$=1, $\alpha$=48 8.219 0.527 0.020 78.70 65.25* 54.98* 65.51* 71.40*
+$\beta$=0.5, $\alpha$=8 13.775 6.082 0.018 80.83 62.49 63.13 65.67 69.04
+$\beta$=0.5, $\alpha$=32 20.876 13.184 0.029 75.58 60.89 68.73 71.71 76.41
+$\beta$=0.5, $\alpha$=48 22.069 14.377 0.006 72.40 60.29 69.48 70.64 79.46
+According to the comparison results in Table 7, the watermark texts generated by DAIR have excellent quality and consistency, and it is difficult for unauthorized receivers to detect the existence of the watermark text, avoiding the problem of watermark abuse.
 Only authorized receivers with the same model settings can detect and fully extract the watermark information.
@@
 Secret can be not only texts, but also other modalities.
-In addition, we embed the “AAAI 2025 icon”.
-For details, see Table 14 in Supplementary A.5.
+In addition, we embed the image information.
+For details, see Table 15 in Appendix 7.4.
 Secret&Bitsream Instruction Stegos PPL $\Delta$LDA
@@
 This makes it difficult to maintain logical coherence when generating longer stegos.
-So longer stegos with high quality is a problem, and there is little research in this task.
+So, longer stegos with high quality is a problem, and there is little research in this task.
 Fortunately, we found that DAIRstega also has this ability.
 Examples of embedded text and cross-modal secrets are shown in Table 8 and Table 15.
-3.6 Ablation experiment
-We performed ablation experiment on different LLMs, and the results are shown in Table 8.
+4.6 Ablation of the different LLMs
+We performed ablation experiments on different LLMs, and the results are shown in Table 9.
 Table 9: Ablation on different LLMs. Bold is the best result. Here we only take “$\beta$ = 1, $\alpha$ = 32” as an example.
-$\beta$=1, $\alpha$=32 PPL $\downarrow$ CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$
-LLaMA2-7B 8.184 97.65 38.48 0.217
-LLaMA2-13B 8.431 96.32 43.93 0.271
-$\beta$=1, $\alpha$=32 MD $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$ BLEU $\uparrow$
-LLaMA2-7B 6.293 0.022 76.97 77.85
-LLaMA2-13B 8.034 0.019 88.38 72.31
-$\beta$=1, $\alpha$=32 $\Delta$DP $\downarrow$ Score $\uparrow$ TS_CSW $\downarrow$ UP4LS $\downarrow$
-LLaMA2-7B 2.35 65.22 54.05 71.97
-LLaMA2-13B 3.68 61.66 54.73 71.12
+$\beta$=1, $\alpha$=32 Perceptual Statistical
+PPL $\downarrow$ CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$ MD $\downarrow$ $\Delta$DP $\downarrow$
+LLaMA2-7B 8.184 97.65 38.48 0.217 6.293 2.35
+LLaMA2-13B 8.431 96.32 43.93 0.271 8.034 3.68
+$\beta$=1, $\alpha$=32 Semantic Anti-steganalysis
+Mauve $\uparrow$ BLEU $\uparrow$ $\Delta$LDA $\downarrow$ Score $\uparrow$ TS_CSW $\downarrow$ UP4LS $\downarrow$
+LLaMA2-7B 76.97 77.85 0.022 65.22 54.05 71.97
+LLaMA2-13B 88.38 72.31 0.019 61.66 54.73 71.12
 According to Table 9, it is not possible to generate stegos with stronger concealment by using larger LLMs.
-This is because LS needs to embed secret.
-The stego quality is more relevant with embedding way rather than using larger LLMs.
+This is because steganography needs to embed secrets.
+The stego quality is more relevant with the embedding way rather than using larger LLMs.
 5 Conclusion
-To enhance the concealment of stegos, this paper proposes an LS scheme based on dynamically allocated intervals, namely DAIRstega.
-It uses the idea of non-uniform roulette and takes the CP of tokens as an important basis for allocating the number of codings.
-Then the requirements of invisible information hiding are analyzed and three constraints are given.
-The allocation expression is condensed.
+To enhance the concealment of stegos, this paper proposes a steganography scheme based on allocated roulette area, namely DAIRstega.
+It uses the idea of the non-uniform roulette wheel and takes the CP of tokens as an important basis for allocating the number of codings.
+Then, the allocation function and three constraints are designed to optimize the allocation process.
 Experiments show that the proposed embedding way and DAIRstega have achieved excellent performance in various concealment and anti-steganalysis.
-We also prove that the stego quality depends more on the design of the model and embedding way, not just the use of LLMs.
-Furthermore, DAIRstega can be used as a secure watermarking scheme.
-Last but not least, to concisely show the proposed scheme, we did not conduct the results of all functions and LLMs.
+Last but not least, to show the proposed scheme concisely, we did not conduct the results of all allocation functions and LLMs.
 6 Outlook
-We have stimulated thinking in this field and discovered gaps in the field for scholars to conduct follow-up research.
+By reflecting on this field, we found gaps in it that can be studied by scholars.
 1.
-Given the limited studies on longer and more diverse LS.
-The scheme can be designed to improve the length and diversity.
+This paper preliminarily explores the performance of the proposed method in long stegos.
+However, the research on long stegos is still in the basic stage.
+So, given the limited studies on longer and more diverse steganography, the scheme can be designed to improve the length and diversity.
 2.
-We know that fine-tuning LLMs brings huge costs, so the fine-tuning-free LS can be improved.
-Knowledge Graphs or Chain of Thought are used to enhance the prompts, and encoding text features to achieve a convenient and completely concealed LS scheme.
+There is currently a lack of high-quality steganography schemes for closed-source LLMs.
+So, knowledge Graphs or Chain of Thought can enhance the prompts, while encoding text features allows for convenient and completely concealed steganography.
 3.
-For watermarking, the idea of LS design can be used to take the hyperparameters of the open-source LLMs as the part that needs to be verified, and construct a new watermarking scheme.
+The scheme in this paper can also be used as a watermarking for LLMs, which can make up for the insufficiency of the watermarking for LLMs being difficult to completely extract.
+Therefore, a new watermarking for LLMs can be designed by referring to the scheme in this paper.
 4.
-Virtual social robots can be built, leveraging secrets to generate multi-turn conversations.
-It promises contactless covert communication, keeping the sender safe.
+At present, linguistic steganography can achieve strong concealment and anti-steganalysis capabilities, that is, it is relatively safe.
+However, the sender of the stego is dangerous.
+Therefore, virtual social robots can be developed that utilize secrets to generate multi-turn conversations, promising contactless, covert communication that ensures the sender’s safety.
 5.
-We recognize the shortcomings of current steganalysis for detecting DAIRstega’s stegos.
-So steganalysis needs to be developed.
-Appendix A Supplementary
-A.1 Related work
-The concealment of stego determines the success of covert communication to a great extent. The concealment is primarily manifested in three aspects: perceptual, statistical, and semantic.
-Perceptual concealment” focuses on ensuring that the scheme generates stego with complete and fluent.
-This is the most basic requirement of LS.
-In pursuit of this objective, Fang et al.
-(Fang, Jaggi, and Argyraki 2017) proposed an LSTM-based LS scheme.
-This scheme segments the vocabulary into several sets based on bit blocks.
-It then selects tokens with the highest probability that corresponds to the secret information from the candidate pool.
-Ding et al.
-(Ding et al.
-2023a) combined the conditional generation strategy with the replacement technique, using text sequences as auxiliary data in the stego generation process to enhance the embedding capabilities.
-Yang et al.
-(Yang et al.
-2019) trained a language model to generate stegos.
-Using HC-based (Huffman coding) and FLC-based (Fixed length coding) to encode the candidate pool, these stegos have excellent completeness and fluentness.
-Statistical concealment” requires the stegos distribution to be close to the covers.
-To achieve this goal, Yang et al.
-(Yang et al.
-2021) designed an encoder-decoder structure, and VLC-based (Variable length coding) is used to encode the CP.
-To further reduce the distribution difference between cover and stego, Zhang et al.
-(Zhang et al.
-2021) used adaptive dynamic grouping coding to recursively embed secret.
-Zhou et al.
-(Zhou et al.
-2021) used a GAN (Generative Adversarial Network) to design an adaptive probability distribution steganography.
-These ensure statistical concealment.
-Semantic concealment” aims to generate the stego that is coherent and semantically controllable.
-To this end, Li et al.
-(Li et al.
-2021) put forward an LS scheme based on the knowledge graph.
-Yang et al.
-(Yang et al.
-2023) utilized semantic information encoding to embed secret information, realizing the effect of maintaining semantics and increasing the embedding capacity during the translation process.
-Wang et al.
-(Wang et al.
-2023a) leveraged the relevance of social network context to enhance contextual semantic relevance while maintaining embedding rates.
-Lu et al.
-(Lu et al.
-2023) used contextual learning and GPT2 guidance to generate stegos.
-A.2 Supplement the effect of the proposed embedding way
-Table 9: Text quality, statistical analysis, and discourse-matching comparison of stegos generated by DAIRstega and baselines.
-The embedding rates of each way include 1 Bpw (bit per word) and 2 Bpw.
-“Avg.” represents the overall performance.
-Bold is the best overall result.
-“ *” is the suboptimal overall result.
-The meanings of “bit” are the same in Table 2.
-Embedding Param / Bpw Text quality Statistical analysis Discourse-matching
-PPL $\downarrow$ $\Delta$Pcs $\downarrow$ CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$ MD $\downarrow$ $\Delta$DP $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$ BLEU $\uparrow$ Score $\uparrow$
-LLMs+HC bit = 1 / 1.00 13.331 5.639 96.54 45.79 0.263 8.255 3.46 0.014 70.60$\pm$27.36 54.26$\pm$4.25 58.44$\pm$4.07
-bit = 2 / 2.52 24.334 16.641 90.57 47.35 0.350 9.354 9.43 0.009 64.88$\pm$30.03 55.00$\pm$5.77 56.19$\pm$4.01
-Avg. 18.832 11.140 93.55 46.57 0.307 8.805 6.45 0.011 67.74 54.63 57.32
-LLMs+ADG $\tau$ = 0.3 / 1.34 12.750 5.058 97.20 45.59 0.240 8.316 2.80 0.010 77.86$\pm$23.65 71.79$\pm$3.03 59.18$\pm$5.84
-$\tau$ = 0.65 / 2.99 16.175 8.483 95.95 44.74 0.331 8.478 4.05 0.018 75.60$\pm$27.02 62.60$\pm$4.68 58.60$\pm$5.69
-Avg. 14.463* 6.770* 96.58 45.17* 0.285 8.397* 3.42 0.014 76.73* 67.20* 58.35*
-LLMs+AC CP = 8 / 1.25 73.454 65.762 82.16 51.64 0.746 15.415 17.84 0.014 35.64$\pm$22.65 50.92$\pm$29.30 45.10$\pm$3.07
-CP = 32 / 1.32 86.238 78.545 80.38 51.21 0.829 14.194 19.62 0.012 37.18$\pm$24.59 48.01$\pm$27.45 47.51$\pm$2.76
-Avg. 79.846 72.154 81.27 51.43 0.788 14.805 18.73 0.013* 36.41 49.47 46.31
-LLMs+FLC bit = 1 / 1.34 12.883 5.190 97.04 46.56 0.241 8.280 2.96 0.031 79.48$\pm$26.53 54.88$\pm$6.33 58.44$\pm$4.64
-bit = 2 / 2.67 30.481 22.788 93.86 46.30 0.328 8.939 6.14 0.011 61.10$\pm$29.30 58.21$\pm$9.88 55.20$\pm$3.39
-Avg. 21.682 13.989 95.45 46.43 0.284* 8.610 4.55 0.021 70.29 56.55 56.82
-Ours $\beta$=1 $\alpha$=8 / 1.10 7.686 0.006 97.20 38.86 0.237 6.458 2.80 0.017 79.91$\pm$26.75 77.88$\pm$30.40 65.76$\pm$4.13
-$\beta$=1 $\alpha$=32 / 1.11 8.184 0.492 97.65 38.48 0.217 6.293 2.35 0.022 76.97$\pm$26.47 77.85$\pm$30.42 65.22$\pm$4.61
-$\beta$=1 $\alpha$=48 / 1.13 8.219 0.527 97.45 38.46 0.226 6.273 2.55 0.020 78.70$\pm$25.51 76.53$\pm$31.71 65.25$\pm$4.78
-Avg. (1Bpw) 8.030 0.342 97.43 38.60 0.227 6.341 2.57 0.020 78.53 77.42 65.41
-[…]
-$\beta$=0.5 $\alpha$=32 / 2.56 20.876 13.184 92.43 39.69 0.389 8.976 7.57 0.029 75.58$\pm$20.11 66.99$\pm$27.08 60.89$\pm$3.85
-$\beta$=0.5 $\alpha$=48 / 2.58 22.069 14.377 92.44 38.31 0.389 8.592 7.56 0.006 72.40$\pm$26.33 79.64$\pm$12.47 60.29$\pm$4.00
-Avg. (2Bpw) 18.907 11.214 94.09 38.60 0.336 8.258 5.91 0.018 76.27 73.27 61.22
-[…]
-Table 10: Anti-steganalysis comparison of stegos generated by DAIRstega and baselines.
-The embedding rates of each way include 1Bpw (bit per word) and 2Bpw.
-“Avg.” represents the overall performance.
-Bold is the best overall result.
-“ *” is the suboptimal overall result.
-Embedding Param / Bpw LS_CNN TS_CSW EILG UP4LS
-Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$
-LLMs+HC bit = 1 / 1.00 77.62$\pm$1.38 78.41$\pm$1.65 78.09$\pm$1.33 78.25$\pm$1.21 78.58$\pm$1.69 78.29$\pm$2.24 86.73$\pm$1.32 86.13$\pm$0.50
-bit = 2 / 2.52 77.54$\pm$1.50 77.57$\pm$2.05 76.03$\pm$1.74 76.09$\pm$1.33 78.29$\pm$0.72 78.95$\pm$1.47 88.19$\pm$1.00 88.43$\pm$2.61
-Avg. 77.58 77.99 77.06 77.17 78.44 78.62 87.46 87.28
-LLMs+ADG $\tau$ = 0.3 / 1.34 70.10$\pm$0.66 70.17$\pm$0.89 69.45$\pm$2.05 69.34$\pm$2.53 70.99$\pm$3.15 70.49$\pm$1.11 84.00$\pm$0.83 84.23$\pm$1.07
-$\tau$ = 0.65 / 2.99 63.06$\pm$2.49 63.69$\pm$3.61 62.22$\pm$1.46 61.86$\pm$1.18 66.14$\pm$1.38 66.15$\pm$1.38 80.42$\pm$0.59 80.57$\pm$0.58
-Avg. 66.58* 66.93* 65.84* 65.60* 68.57* 68.32* 82.21* 82.40*
-LLMs+AC CP = 8 / 1.25 90.15$\pm$1.17 89.98$\pm$1.21 89.69$\pm$0.93 89.37$\pm$0.88 86.92$\pm$1.12 86.52$\pm$0.96 93.62$\pm$0.89 93.37$\pm$0.40
-CP = 32 / 1.32 88.45$\pm$1.38 88.44$\pm$1.25 90.80$\pm$1.97 90.52$\pm$2.18 86.44$\pm$2.47 85.86$\pm$3.10 94.03$\pm$0.64 94.16$\pm$0.95
-Avg. 89.30 89.21 90.25 89.95 86.68 86.19 93.83 93.77
-LLMs+FLC bit = 1 / 1.34 78.53$\pm$1.20 80.05$\pm$1.39 82.72$\pm$1.27 83.32$\pm$1.10 76.67$\pm$0.85 76.15$\pm$1.96 89.03$\pm$0.25 89.20$\pm$0.80
-bit = 2 / 2.67 76.49$\pm$1.15 76.54$\pm$1.13 77.08$\pm$2.06 77.26$\pm$1.75 76.95$\pm$1.69 76.46$\pm$1.74 90.62$\pm$0.89 90.37$\pm$0.40
-Avg. 77.51 78.30 79.90 80.29 76.81 76.31 89.83 89.79
-Ours $\beta$=1 $\alpha$=8 / 1.10 59.50$\pm$2.77 60.15$\pm$10.59 56.82$\pm$4.09 64.21$\pm$10.03 67.25$\pm$3.23 68.04$\pm$1.53 75.40$\pm$0.37 76.26$\pm$1.28
-$\beta$=1 $\alpha$=32 / 1.11 59.35$\pm$1.25 59.99$\pm$5.51 54.05$\pm$4.72 61.68$\pm$11.17 62.28$\pm$0.50 62.66$\pm$2.31 71.97$\pm$0.64 72.75$\pm$1.21
-$\beta$=1 $\alpha$=48 / 1.13 58.00$\pm$2.37 56.08$\pm$13.79 54.98$\pm$3.14 59.17$\pm$11.37 65.51$\pm$2.23 66.78$\pm$3.91 71.40$\pm$0.65 70.89$\pm$2.05
-Avg. (1Bpw) 58.95 58.74 55.28 61.69 65.01 65.83 72.92 73.30
-[…]
-$\beta$=0.5 $\alpha$=32 / 2.56 72.06$\pm$1.17 71.78$\pm$1.21 68.73$\pm$1.29 68.19$\pm$3.44 71.71$\pm$1.95 70.34$\pm$2.41 76.41$\pm$2.56 75.90$\pm$2.73
-$\beta$=0.5 $\alpha$=48 / 2.58 70.17$\pm$1.81 70.40$\pm$3.11 69.48$\pm$0.68 67.80$\pm$1.98 70.64$\pm$1.61 71.00$\pm$2.23 79.46$\pm$0.56 78.28$\pm$1.13
-Avg. (2Bpw) 68.37 68.13 67.11 65.25 69.34 69.06 74.97 74.14
-[…]
-A.3 Complete data for Table 4 and Table 5
-In this section, we give the complete data of Table 4 and Table 5, as shown in Table 11 and Table 12.
-Table 11 contains the standard deviations of many datasets with different discourse, and Table 12 contains the standard deviations of multiple runs.
-Table 11: Complete data for Table 4.
-The meanings of “bin” and “bit” are the same in Table 2.
-“a±$\pm$±b” represents “Avg.$\pm$±Std”.
-Bold is the best result.
-“ *” is the suboptimal result.
-The “Std”s of DAIRstega Mauve and BLEU values are bigger.
-This is not due to the instability of 5-time runs, but is obtained from the effect on dozens of different discourse datasets, and represents the overall performance.
-Schemes Param / Bpw Mauve $\uparrow$ BLEU $\uparrow$ Score $\uparrow$
-Fang bin = 1 / 1.00 2.15$\pm$0.87 0.90$\pm$1.58 42.58$\pm$1.30
-bin = 3 / 3.00 1.96$\pm$0.68 0.90$\pm$1.59 41.65$\pm$1.33
-RNNstega bit = 1 / 1.00 12.73$\pm$10.56 1.87$\pm$3.28 52.57$\pm$2.20
-bit = 3 / 2.63 11.97$\pm$10.56 1.83$\pm$3.21 52.36$\pm$1.80
-ADG $\tau$ = 0.5 / 4.38 18.54$\pm$13.89 2.72$\pm$4.79 47.55$\pm$2.25
-LSCS - / 1.12 2.40$\pm$1.78 31.95$\pm$12.89 59.22$\pm$3.14
-PLMmark - / - 7.22$\pm$4.88 0.15$\pm$0.27 43.50$\pm$0.79
-DAIRstega $\beta$ = 1 $\alpha$ = 8 / 1.10 79.91$\pm$26.75 76.37$\pm$31.85 65.76$\pm$4.13
-$\alpha$ = 32 / 1.11 76.97$\pm$26.47 77.85$\pm$30.42* 65.22$\pm$4.61
-$\alpha$ = 48 / 1.13 78.70$\pm$25.51 76.53$\pm$31.71 65.25$\pm$4.78*
-[…]
-$\alpha$ = 32 / 2.56 75.58$\pm$20.11 66.99$\pm$27.08 60.89$\pm$3.85
-$\alpha$ = 48 / 2.58 72.40$\pm$26.33 79.64$\pm$12.47 60.29$\pm$4.00
-Table 12: Complete data for Table 5.
-The meanings of “bin” and “bit”are the same in Table 2.
-“a±$\pm$±b” represents “Avg.$\pm$±Std”.
-“Avg” and “Std” in this table are obtained by running 5 times.
-Bold is the best result.
-“ *” is the suboptimal result.
-Schemes Param / Bpw LS_CNN TS_CSW EILG UP4LS
-Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$ Acc $\downarrow$ F1 $\downarrow$
-Fang bin = 1 / 1.00 99.70$\pm$0.29 99.68$\pm$0.31 88.34$\pm$21.22 91.01$\pm$15.69 99.50$\pm$0.25 99.51$\pm$0.25 99.75$\pm$0.06 99.60$\pm$0.09
-bin = 3 / 3.00 99.75$\pm$0.16 99.76$\pm$0.15 89.03$\pm$20.58 92.39$\pm$13.82 99.59$\pm$0.17 99.52$\pm$0.66 99.88$\pm$0.10 99.80$\pm$0.15
-RNNstega bit = 1 / 1.00 85.81$\pm$1.93 86.03$\pm$1.40 81.39$\pm$1.70 82.59$\pm$1.62 85.86$\pm$0.50 86.29$\pm$0.89 97.42$\pm$0.44 95.97$\pm$0.67
-bit = 3 / 2.63 82.93$\pm$2.18 84.48$\pm$2.20 80.35$\pm$2.38 77.94$\pm$3.04 82.71$\pm$0.17 83.10$\pm$1.02 96.78$\pm$0.82 94.74$\pm$1.34
-ADG $\tau$ = 0.5 / 4.38 85.86$\pm$1.73 86.35$\pm$1.82 82.23$\pm$2.40 82.08$\pm$2.96 84.86$\pm$1.74 84.10$\pm$1.80 98.34$\pm$0.39 97.41$\pm$0.58
-LSCS - / 1.12 78.86$\pm$3.07 78.98$\pm$2.83 81.19$\pm$6.62 80.11$\pm$8.26 84.15$\pm$0.41 85.11$\pm$0.79 96.22$\pm$2.08 95.48$\pm$3.11
-PLMmark - / - 98.64$\pm$0.92 98.60$\pm$0.95 98.81$\pm$1.83 98.75$\pm$1.93 99.92$\pm$0.17 99.92$\pm$0.11 99.85$\pm$0.30 99.85$\pm$0.30
-DAIRstega $\beta$ = 1 $\alpha$ = 8 / 1.10 59.50$\pm$2.77 60.15$\pm$10.59 56.82$\pm$4.09 64.21$\pm$10.03 67.25$\pm$3.23 68.04$\pm$1.53 75.40$\pm$0.37 76.26$\pm$1.28
-$\alpha$ = 32 / 1.11 59.35$\pm$1.25 59.99$\pm$5.51* 54.05$\pm$4.72 61.68$\pm$11.17 62.28$\pm$0.50 62.66$\pm$2.31* 71.97$\pm$0.64 72.75$\pm$1.21
-$\alpha$ = 48 / 1.13 58.00$\pm$2.37 56.08$\pm$13.79 54.98$\pm$3.14* 59.17$\pm$11.37 65.51$\pm$2.23 66.78$\pm$3.91 71.40$\pm$0.65* 70.89$\pm$2.05*
-[…]
-$\alpha$ = 32 / 2.56 72.06$\pm$1.17 71.78$\pm$1.21 68.73$\pm$1.29 68.19$\pm$3.44 71.71$\pm$1.95 70.34$\pm$2.41 76.41$\pm$2.56 75.90$\pm$2.73
-$\alpha$ = 48 / 2.58 70.17$\pm$1.81 70.40$\pm$3.11 69.48$\pm$0.68 67.80$\pm$1.98 70.64$\pm$1.61 71.00$\pm$2.23 79.46$\pm$0.56 78.28$\pm$1.13
-A.4 Supplement data for anti-steganalysis
-We also perform other 7 steganalysis methods to detect the stegos we generated, as shown in Table A.4. These include non-BERT-based, BERT-based, and LLMs-based steganalysis methods.
-Table 13: The supplemental anti-steganalysis capability of the DAIRstega from other 7 steganalysis works.
-Bold represents the best result and the best Avg.
-result.
-“ *” represents the suboptimal result.
+Finally, we found that existing steganalysis methods have poor detection performance on DAIRstega’s stegos.
+Therefore, it is essential to advance steganalysis techniques.
+Acknowledgment
+We would like to express our gratitude to the MindSpore Community for their invaluable support and resources throughout this research.
+Thanks for the support provided by MindSpore Community.
+Thank the editor and reviewers of “Applied Soft Computing” for their review and constructive suggestions.
+Declaration of competing interest
+The authors declare no conflict of interest.
+Funding
+This work is supported by the National Natural Science Foundation of China (Grant U21B2020) and supported by BUPT Excellent Ph.D. Students Foundation (Grant CX2023120, CX20241055).
+Author contributions
+Yihao Wang: Conceptualization; Data curation; Investigation; Methodology; Software; Funding acquisition; original draft.
+Ruiqi Song: Data curation; Investigation; Resources; Software; Validation; original draft.
+Lingxiao Li: Data curation; Resources; Software; Funding acquisition; original draft; review & editing.
+Yihao Wang and Lingxiao Li contributed equally to this work and should be considered co-first authors.
+Ru Zhang: Funding acquisition; Investigation; Project administration; Supervision; review & editing.
+Jianyi Liu: Formal analysis; Supervision; Investigation; Validation; review & editing.
+Ethical Statement
+This research does not involve human participants, personal data, or animal subjects; therefore, it does not raise any direct ethical concerns.
+The data used in this study were sourced from publicly available datasets or generated by applicable laws and regulations.
 Supplemental VAE-Stega [3] DAIRstega
@@
 F1 $\downarrow$ 92.19 73.20 95.62 94.53 88.26 97.60 90.23 72.31 66.19 66.29* 71.47 70.02 69.04 69.22
-LLMs- [7] Acc $\downarrow$ 89.63 78.13 93.13 88.63 76.38 86.75 85.44 63.38 59.63* 61.44 63.25 59.38 62.13 61.53
+LLM- [38] Acc $\downarrow$ 89.63 78.13 93.13 88.63 76.38 86.75 85.44 63.38 59.63* 61.44 63.25 59.38 62.13 61.53
 F1 $\downarrow$ 89.59 78.06 93.12 88.62 76.13 86.75 85.38 63.24 60.15 60.59* 63.13 59.15 61.86 61.35
@@
 F1 $\downarrow$ 96.87 92.97 98.57 97.10 92.97 98.97 96.25 75.27 72.41 71.35* 81.55 69.32 83.15 75.51
-  • [1] VAE-Stega: Linguistic Steganography Based on Variational Auto-Encoder. IEEE Trans. Inf. Forensics and Security, 880–895, 2021.
-  • [2] A Fast and Efficient Text Steganalysis Method. IEEE Signal Processing Lett., 26(4): 627–631, 2019.
-  • [3] A Hybrid R-BiLSTM-C Neural Network Based Text Steganalysis. IEEE Signal Processing Lett., 26(12): 1907–1911, 2019.
-  • [4] High-performance Linguistic Steganalysis, Capacity Estimation and Steganographic Positioning. In Proceeding of IWDW, 2021.
-  • [5] Small-Scale Linguistic Steganalysis for Multi-Concealed Scenarios. IEEE Signal Processing Lett., 29: 130–134, 2022.
-  • [6] SeSy: Linguistic Steganalysis Framework Integrating Semantic and Syntactic Features. IEEE Signal Processing Lett., 31–35, 2021.
-  • [7] Towards Next-Generation Steganalysis: LLMs Unleash the Power of Detecting Steganography. arXiv:2405.09090, 2024.
-  • [8] Linguistic Steganalysis via LLMs: Two Modes for Efficient Detection of Strongly Concealed Stego. arXiv:2406.04218, 2024.
-A.5 Example of cross-modal secret information hiding
-Table 14: Examples of longer stegos generated by DAIRstega using cross-modal information. Bitstream is the binary corresponding to secret.
+7.4 Example of cross-modal secret information hiding
+Table 15: Examples of longer stegos generated by DAIRstega using cross-modal information. Bitstream is the binary corresponding to the secret.
 Secret&Bitsream Instruction Stegos PPL $\Delta$LDA
@@
 causes of the
-Great Depression.
+Great Depre-
+ssion.
 The Great Depression was a severe worldwide economic depression that took place from 1929 to 1939.
@@
 The collapse of agricultural prices resulted from overproduction, due to improvements in seed stock, fertilizers, pesticides, transportation, and farm equipment, as well as increases in acreage and livestock numbers.
-2.5611
+2.561
 0.021
@@
 Finally, in the field of weather forecasting, quantum computers can be used to simulate the complex interactions between different weather systems, which could lead to more accurate and precise forecasts.
-1.9153
+1.915
 0.009